### v1

# The iterated Dirichlet process and applications to Bayesian inference

Abstract
Consider an i.i.d. sequence of random variables, taking values in some space $S$, whose underlying distribution is unknown. In problems of Bayesian inference, one models this unknown distribution as a random measure, and the law of this random measure is the prior. When $S=\{0,1\}$, a commonly used prior is the uniform distribution on $[0,1]$, or more generally, the beta distribution. When $S$ is finite, the analogous choice is the Dirichlet distribution. For a general space $S$, we are led naturally to the Dirichlet process (see Ferguson [5]).
Here, we consider an array of random variables, and in so doing are led to what we call the iterated Dirichlet process (IDP). We define the IDP and then show how to compute the posterior distribution, given a finite set of observations, using the method of sequential imputation. Ordinarily, this method requires the existence of certain joint density functions, which the IDP lacks. We therefore present a new, more general proof of the validity of sequential imputation, and show that the hypotheses of our proof are satisfied by the IDP.
AMS subject classifications: Primary 62G05; secondary 60G57, 62D10, 62M20
Keywords and phrases: exchangeability, Dirichlet processes, Bayesian inference, importance sampling, sequential imputation
1 Introduction
Let $S$ be a complete and separable metric space and $\{\eta_{i}:i\in\mathbb{N}\}$ an exchangeable sequence of $S$-valued random variables. By de Finetti’s theorem, there exists a random measure $\lambda$ on $S$ such that, given $\lambda$, the sequence $\eta_{1},\eta_{2},\ldots$ is conditionally i.i.d. with distribution $\lambda$. To compute the conditional distribution of $\eta_{N+1},\eta_{N+2},\ldots$ given $\eta_{1},\ldots,\eta_{N}$, we must, among other things, assign a prior distribution to $\lambda$.
The random measure $\lambda$ is an element of $M_{1}=M_{1}(S)$, the space of probability measures on $S$. A prior distribution for $\lambda$ is therefore an element of $M_{1}(M_{1}(S))$. In the case $S=\{0,1\}$, we can identify a measure $\nu\in M_{1}(S)$ with its projected value $\nu(\{1\})\in[0,1]$. In this way, we can identify $M_{1}(M_{1}(S))$ with $M_{1}([0,1])$. Hence, a prior distribution for $\lambda$ is simply a probability measure on $[0,1]$. A common choice for a non-informative prior in this case is the uniform distribution, or its generalization, the beta distribution.
In the case $S=\{0,\ldots,d\}$, we can identify a measure $\nu\in M_{1}(S)$ with its projected values $(\nu(\{0\}),\ldots,\nu(\{d\}))\in\Delta^{d}$, where $\Delta^{d}$ is the standard $d$-simplex,
${\Delta^{d}=\{t=(t_{0},\ldots,t_{d})\in\mathbb{R}^{d+1}:t_{j}\geq 0\text{ and }\sum_{0}^{d}t_{j}=1\}.}$ (1.1)
In this way, a prior distribution for $\lambda$ is simply a probability measure on $\Delta^{d}$. A common choice in this case is the Dirichlet distribution.
In the case of a general $S$, a common choice for a non-informative, non-parametric prior is to let $\lambda$ be a Dirichlet process on $S$. The Dirichlet process was first introduced by Thomas S. Ferguson in [5]. Its definition and relevant properties are formally summarized in Section 3. A Dirichlet process is almost surely a discrete measure. It has two parameters, $\kappa$ and $\rho$. The probability measure $\rho\in M_{1}(S)$ is its mean measure, or base measure, satisfying $E[\lambda(A)]=\rho(A)$. The number $\kappa\in(0,\infty)$ is its concentration parameter. The smaller $\kappa$ is, the more likely $\lambda$ is to be concentrated on only a few points in $S$. If $\lambda$ is a Dirichlet process on $S$ with parameters $\kappa$ and $\rho$, then its law, which is an element of $M_{1}(M_{1}(S))$, is denoted by $\mathcal{D}(\kappa\rho)$.
Now let $\{\xi_{ij}:i,j\in\mathbb{N}\}$ be an array of $S$-valued random variables. In [4], such arrays were intended to model a situation in which there are several agents, all from the same population, each undertaking a sequence of actions. The space $S$ represents the set of possible actions and $\xi_{ij}$ represents the $j$th action of the $i$th agent. A simple example presented in [4] was the following. Imagine a pressed penny machine, like those found in museums or tourist attractions. For a fee, the machine presses a penny into a commemorative souvenir. Now imagine the machine is broken, so that it mangles all the pennies we feed it. Each pressed penny it creates is mangled in its own way, and has its own probability of landing on heads when flipped. The machine, though, might have certain tendencies. For instance, it might tend to produce pennies that are biased toward heads. In this situation, the agents are the pennies and the actions are the heads and tails that they produce. In this example, then, we have $S=\{0,1\}$ and the random variable $\xi_{ij}\in S$ denotes the result of the $j$th flip of the $i$th penny created by the machine.
We assume that $\xi$ is row exchangeable, which means that
  1. (i)
    each row of the array, $\xi_{i}=\{\xi_{ij}:j\in\mathbb{N}\}$, is an exchangeable sequence of $S$-valued random variables, and
  2. (ii)
    the sequence of rows, $\{\xi_{i}:i\in\mathbb{N}\}$, is an exchangeable sequence of $S^{\infty}$-valued random variables.
It is easy to see that $\xi$ is row exchangeable if and only if $\{\xi_{\sigma(i),\tau_{i}(j)}\}$ and $\{\xi_{ij}\}$ have the same finite-dimensional distributions whenever $\sigma$ and $\tau_{i}$ are (finite) permutations of $\mathbb{N}$. As a matter of notation, we will sometimes write $\xi(i)$ and $\xi(i,j)$ for $\xi_{i}$ and $\xi_{ij}$, respectively.
According to a de Finetti theorem for row exchangeable arrays (see [4, Theorem 3.2]), there exists a sequence of random measure $\mu_{1},\mu_{2},\ldots$ on $S$ and a random measure $\varpi$ on $M_{1}(S)$ such that
  • given $\varpi$, the sequence $\mu_{1},\mu_{2},\ldots$ is i.i.d. with distribution $\varpi$, and
  • for each $i$, given $\mu_{i}$, the sequence $\xi_{i1},\xi_{i2},\ldots$ is i.i.d. with distribution $\mu_{i}$.
The random measures $\mu_{i}$ are called the row distributions of $\xi$, and $\varpi$ is called the row distribution generator.
In the present work, we are concerned with performing Bayesian inference on row exchangeable arrays. The first step toward this to assign a prior distribution to $\xi$. According to the above, this is equivalent to assigning to a prior distribution to $\varpi$. Following Ferguson, we would like to define $\varpi$ to be a Dirichlet process. Since $\varpi$ is a random measure on $M_{1}(S)$, the law of $\varpi$ should have the form $\mathcal{D}(\kappa\rho)$, where $\kappa>0$ and $\rho$ is a nonrandom measure on $M_{1}(S)$. If $B\subseteq M_{1}(S)$ is measurable, then
$P(\mu_{i}\in B)=E[P(\mu_{i}\in B\mid\varpi)]=E[\varpi(B)]=\rho(B).$
In other words, $\rho$ is the prior distribution for $\mu_{i}$, which is a random measure on $S$. Again turning to the Dirichlet process, we let $\rho=\mathcal{D}(\varepsilon\varrho)$, where $\varepsilon>0$ and $\varrho$ is a nonrandom measure on $S$. As above, if $A\subseteq S$ is measurable, then
$P(\xi_{ij}\in A)=E[P(\xi_{ij}\in A\mid\mu_{i})]=E[\mu_{i}(A)]=\varrho(A),$
so that $\varrho$ is the prior distribution for $\xi_{ij}$. This means that $\varpi$ has distribution $\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$. We call a random measure with a distribution of this form an iterated Dirichlet process (or IDP) on $S$ with parameters $\kappa$, $\varepsilon$, and $\varrho$. We call $\kappa$ and $\varepsilon$ the column and row concentrations, respectively, and we call $\varrho$ the base measure of $\varpi$. The effect of $\kappa$ and $\varepsilon$ on the behavior of $\varpi$ is discussed in Section 3.5.
Having assigned a prior, we turn our attention to inference. We adopt the notation used in [4]:
$X_{in}=\begin{pmatrix}\xi_{i1}&\cdots&\xi_{in}\end{pmatrix},\qquad{\bm{X}}_{mn}=\begin{pmatrix}X_{1n}\\ \vdots\\ X_{mn}\end{pmatrix}=\begin{pmatrix}\xi_{11}&\cdots&\xi_{1n}\\ \vdots&\ddots&\vdots\\ \xi_{m1}&\cdots&\xi_{mm}\end{pmatrix},\qquad\bm{\mu}_{m}=\begin{pmatrix}\mu_{1}\\ \vdots\\ \mu_{m}\end{pmatrix}.$
Our objective is to make inferences about the future of the process $\xi$ based on past observations. That is, if $M,N,N^{\prime}\in\mathbb{N}$ and $N<N^{\prime}$, then we wish to compute
$\mathcal{L}(\xi_{ij}:i\leq M,\;N<j\leq N^{\prime}\mid{\bm{X}}_{MN}).$ (1.2)
Here, the notation $\mathcal{L}(X\mid Y)$ denotes the regular conditional distribution of $X$ given $Y$.
The exact computation of (1.2) is intractable. For example, consider the simplest nontrivial case, $M=2$ and $S=\{0,1\}$. When $S=\{0,1\}$, an IDP on $S$ can be regarded as simply a Dirichlet process on $[0,1]$ whose base measure is a Beta distribution. Such a process is considered in [3]. There, the authors provide us with a formula for
$P(\xi_{2,N+1}=1\mid\xi_{ij}=k_{ij},i=1,2,j=1,\ldots,N).$
To state this formula, let $\varrho_{k}=\varrho(\{k\})$, $\alpha_{k}=\varepsilon\varrho_{k}$, and $y_{i}=\sum_{j=1}^{N}k_{ij}$. Using the notation $t^{(n)}=\Gamma(t+n)/\Gamma(t)=\prod_{k=0}^{n-1}(t+k)$, define
$r={\frac{\kappa\varepsilon^{(2N)}}{(\varepsilon^{(N)})^{2}}\frac{\alpha_{1}^{(y_{1})}\alpha_{1}^{(y_{2})}}{\alpha_{1}^{(y_{1}+y_{2})}}\frac{\alpha_{0}^{(N-y_{1})}\alpha_{0}^{(N-y_{2})}}{\alpha_{0}^{(2N-y_{1}-y_{2})}}}.$
We then have
$P(\xi_{2,N+1}=1\mid\xi_{ij}=k_{ij},i=1,2,j=1,\ldots,N)\\ ={\frac{r}{1+r}\left(\frac{\varepsilon}{\varepsilon+N}\varrho_{1}+\frac{N}{\varepsilon+n}\frac{y_{2}}{N}\right)+\frac{1}{1+r}\left(\frac{\varepsilon}{\varepsilon+2N}\varrho_{1}+\frac{2N}{\varepsilon+2N}\frac{y_{1}+y_{2}}{2N}\right)}.$
The case $M=3$ and $S=\{0,1\}$ is also outlined in [3], and involves roughly twice as many computations. To deal with larger values of $M$, as well as more general spaces $S$, we need to adopt a different approach to the computation of (1.2).
In [4], we showed that (1.2) can be computed in terms of the distributions,
$\mathcal{L}(\mu_{m}\mid X_{mN},\ldots,X_{MN},\bm{\mu}_{m-1}),$ (1.3)
where $1\leq m\leq M$ (see Section 2.4). But the same issue that makes (1.2) intractable also hinders the computation of (1.3). Namely, in both cases, we are trying to condition on more than one row of the array $\xi$. If, instead of (1.3), we wanted to compute
$\mathcal{L}(\mu_{m}\mid X_{mN},\bm{\mu}_{m-1}),$ (1.4)
for $1\leq m\leq M$, then this can be done (see Section 5.3). The problem, then, is to find a way to use (1.4) to determine (1.2).
This was done by Liu in [12] for the special case $S=\{0,1\}$, using a method known as sequential imputation. The proofs in [12] cite [11], in which Kong, Liu, and Wong first developed the method of sequential imputation. In [11, Section 2], the authors give their core argument, demonstrating the validity of the method. This argument evidently presupposes that the underlying random variables have densities with respect to some product measure. In Proposition 4.5, we show that this is not true for the IDP, not even in the case $S=\{0,1\}$ that was treated by Liu in [12]. In other words, the work in [11] is not sufficient to justify its use in [12].
In Theorem 4.7, we provide a new proof that justifies sequential imputation. Our proof avoids the supposition of a full joint density. In doing so, it shows that sequential imputation can be validly applied to an IDP on any complete and separable metric space. This includes the special case $S=\{0,1\}$ treated in [12] as well as the general case treated here.
When we apply sequential imputation to the IDP, we arrive at our main result, Theorem 5.2. Combined with [4, Theorem 3.4], which is presented below as Theorem 2.4, this allows us to compute (1.2).
The outline of this paper is as follows. In Section 2, we present some basic background results and establish notation. We also include the main results from [4] on row exchangeable arrays. In Section 3, we give the necessary background on Dirichlet processes, including Ferguson’s definition, the Sethuraman stick-breaking construction, mixtures of Dirichlet processes, and the formulas of inference for both ordinary observations as well as noisy observations.
In Section 3.5, we take a closer look at the iterated Dirichlet process, which we defined above. Namely, we discuss the relationship between the parameters, $\kappa$, $\varepsilon$, and $\varrho$, and the behavior of $\varpi$. It should be noted that the iterated Dirichlet process is not the same thing as the hierarchical Dirichlet process, a process that appears elsewhere in the literature (see, for example, [15, 16]). The latter is a mixture of Dirichlet processes, whereas the IDP is not. For a more detailed explanation of the differences between the two, see Remark 3.6.
In Section 4, as described above, we discuss the method of sequential imputation. We then apply this in Section 5 to the iterated Dirichlet process, obtaining our main results, which are Theorem 5.2 and Corollary 5.3. Finally, in Section 6, we present several hypothetical examples to illustrate the use of our main results.
2 Background and notation
2.1 Random measures and kernels
If $(X,\mathcal{U})$ is a topological space, then $\mathcal{B}(X)=\sigma(\mathcal{U})$ denotes the Borel $\sigma$-algebra on $X$. We may sometimes denote $\mathcal{B}(X)$ by $\mathcal{B}_{X}$. We use $\mathcal{R}^{d}$ to denote the Borel $\sigma$-algebra on $\mathbb{R}^{d}$ with the Euclidean topology. If $(S,\mathcal{S})$ is a measurable space and $B\in\mathcal{S}$, then $\mathcal{S}|_{B}=\{A\cap B:A\in\mathcal{S}\}=\{A\in\mathcal{S}:A\subseteq B\}$. We let $\mathbb{R}^{*}=[-\infty,\infty]$ denote the extended real line, equipped with the topology generated by the metric $\rho(x,y)=|\tan^{-1}(x)-\tan^{-1}(y)|$. We use $\mathcal{R}^{*}$ to denote $\mathcal{B}(\mathbb{R}^{*})$. Note that =$\mathcal{R}^{*}=\{A\subseteq\mathbb{R}^{*}:A\cap\mathbb{R}\in\mathcal{R}\}$.
[…]
Let $M_{1}=M_{1}(S)$ be the set of all probability measures on $S$. Note that $M_{1}=\{\nu\in M:\nu(S)\in\{1\}\}$. Hence, $M_{1}\in\mathcal{M}$ and we may define $\mathcal{M}_{1}$ by $\mathcal{M}_{1}=\mathcal{M}|_{M_{1}}$. A random probability measure on $S$ is a function $\mu:\Omega\to M_{1}$ that is $(\mathcal{F},\mathcal{M}_{1})$-measurable, or equivalently, a random measure taking values in $M_{1}$.
We equip $M_{1}$ with the Prohorov metric, $\bm{\pi}$, which metrizes weak convergence. Since $S$ is complete and separable, $M_{1}$ is complete and separable under $\bm{\pi}$. It can be shown that $\mathcal{M}_{1}=\mathcal{B}(M_{1})$. (See, for example, [6, Theorem 2.3].)
Unless otherwise specified, we will equip $S^{\infty}$ with the metric ${\bf d}^{\infty}$ defined by
${\bf d}^{\infty}(x,y)=\sum_{n=1}^{\infty}\frac{{\bf d}(x_{n},y_{n})\wedge 1}{2^{n}}.$
Note that $x\to y$ in $S^{\infty}$ if and only if $x_{n}\to y_{n}$ in $S$ for all $n$. In particular, the metric ${\bf d}^{\infty}$ induces the product topology on $S^{\infty}$. Consequently, $\mathcal{S}^{\infty}=\mathcal{B}(S^{\infty})$ and $S^{\infty}$ is separable. In fact, $(S^{\infty},{\bf d}^{\infty})$ is a complete and separable metric space.
Let $(T,\mathcal{T})$ be a measurable space. A function $\mu:T\to M$ is called a kernel from $T$ to $S$ if $\mu(\cdot,B)$ is $(\mathcal{T},\mathcal{R}^{*})$-measurable for all $B\in\mathcal{S}$. By [7, Lemma 1.37], given a function $\mu:T\to M$ and a $\pi$-system $\mathcal{C}$ such that $\mathcal{S}=\sigma(\mathcal{C})$, the following are equivalent:
  1. (i)
    $\mu$ is a kernel.
  2. (ii)
    $\mu(\cdot,B)$ is $(\mathcal{T},\mathcal{R}^{*})$-measurable for all $B\in\mathcal{C}$.
  3. (iii)
    $\mu$ is $(\mathcal{T},\mathcal{M})$-measurable.
A probability kernel is a kernel taking values in $M_{1}$. Note that a random measure on $S$ is just a kernel from $\Omega$ to $S$, and a random probability measure on $S$ is just a probability kernel from $\Omega$ to $S$.
It can be shown that the function $\mu\mapsto\mu^{\infty}$ mapping $M_{1}$ to $M_{1}(S^{\infty})$ is measurable. Therefore, if $\mu$ is a random probability measure on $S$, then $\mu^{\infty}$ is a random probability measure on $S^{\infty}$. In fact, the random variable $\mu^{\infty}$ is $\sigma(\mu)$-measurable.
Let $S^{\prime}$ be another complete and separable metric space. Let $\gamma$ be a probability kernel from $T$ to $S$ and $\gamma^{\prime}$ a probability kernel from $T\times S$ to $S^{\prime}$. (We allow the possibility that $T$ is a singleton, in which case $\gamma$ is a probability measure on $S$ and $\gamma^{\prime}$ is a probability kernel from $S$ to $S^{\prime}$.) We write $\gamma\gamma^{\prime}$ to denote the probability kernel from $T$ to $S\times S^{\prime}$ characterized by
$(\gamma\gamma^{\prime})(y,A\times A^{\prime})=\int_{A}\gamma^{\prime}(y,z,A^{\prime})\,\gamma(y,dz),$
where $y\in T$, $z\in S$, $A\in\mathcal{S}$, and $A^{\prime}\in\mathcal{S}^{\prime}$. In particular, this means
$\int_{S\times S^{\prime}}f(z,z^{\prime})\,(\gamma\gamma^{\prime})(y,dz\,dz^{\prime})=\int_{S}\int_{S^{\prime}}f(z,z^{\prime})\,\gamma^{\prime}(y,z,dz^{\prime})\,\gamma(y,dz).$
As shorthand for this equation, we write
$(\gamma\gamma^{\prime})(y,dz\,dz^{\prime})=\gamma^{\prime}(y,z,dz^{\prime})\,\gamma(y,dz).$
If $T$ is a singleton, then $\gamma\gamma^{\prime}$ is a probability measure and $(\gamma\gamma^{\prime})(dz\,dz^{\prime})=\gamma^{\prime}(z,dz^{\prime})\,\gamma(dz)$.
2.2 Regular conditional distributions
If $X$ is an $S$-valued random variable and $\mu$ is a probability measure on $S$, then we write $X\sim\mu$, as usual, to mean that $X$ has distribution $\mu$. We also use $\mathcal{L}(X)$ to denote the distribution of $X$, so that $X\sim\mu$ and $\mathcal{L}(X)=\mu$ are synonymous. If $\nu$ is a finite, nonzero measure on $S$, then we write $\mathcal{L}(X)\propto\nu$ to mean that $\mathcal{L}(X)=\nu/\nu(S)$. We will sometimes combine this with differential notation. For example, $d\mathcal{L}(X)\propto f(x)\,\nu(dx)$ means that for some constant $c\in(0,\infty)$, the function $cf$ is the Radon-Nikodym derivative of $\mathcal{L}(X)$ with respect to $\nu$. A naked differential indicates Lebesgue measure, so that in the case $S=\mathbb{R}$, the notation $d\mathcal{L}(X)=f(x)\,dx$ means that $f$ is the density of $X$ with respect to Lebesgue measure. If $\mathcal{L}(X)=\mu$ and $B\in\mathcal{S}$, then we use $\mathcal{L}(X;B)$ to denote $\mu(B)$. In this way, we may sometimes write $\mathcal{L}(X;dx)$ instead of $d\mathcal{L}(X)$.
Let $\mathcal{G}\subseteq\mathcal{F}$ be a $\sigma$-algebra. A regular conditional distribution for $X$ given $\mathcal{G}$ is a random probability measure $\mu$ on $S$ such that $P(X\in A\mid\mathcal{G})=\mu(A)$ a.s. for all $A\in\mathcal{S}$. Since $S$ is a complete and separable metric space, regular conditional distributions exist. In fact, such a $\mu$ exists whenever $X$ takes values in a standard Borel space. Moreover, $\mu$ is unique in the sense that if $\mu$ and $\widetilde{\mu}$ are two such random measures, then $\mu=\widetilde{\mu}$ a.s. That is, with probability one, $\mu(B)=\widetilde{\mu}(B)$ for all $B\in\mathcal{S}$. The random measure $\mu$ may be chosen so that it is $\mathcal{G}$-measurable.
We use the notation $\mathcal{L}(X\mid\mathcal{G})$ to refer to the regular conditional distribution of $X$ given $\mathcal{G}$. If $\mu$ is a random probability measure, the notation ${X\mid\mathcal{G}}\sim\mu$ means $\mathcal{L}(X\mid\mathcal{G})=\mu$ a.s. If $\nu$ is a random measure that is a.s. finite and nonzero, then $\mathcal{L}(X\mid\mathcal{G})\propto\nu$ means ${X\mid\mathcal{G}}\sim\nu/\nu(S)$. As with unconditional distributions, we will combine this with differential and semicolon notation, so that, for instance, $\mathcal{L}({X\mid\mathcal{G}};A)=P(X\in A\mid\mathcal{G})$ a.s.
Let $(T,\mathcal{T})$ be a measurable space and $Y$ a $T$-valued random variable. Then there exists a probability kernel $\mu$ from $T$ to $S$ such that $X\mid\sigma(Y)\sim\mu(Y)$. Moreover, $\mu$ is unique in the sense that if $\mu$ and $\widetilde{\mu}$ are two such probability kernels, then $\mu=\widetilde{\mu}$, $\mathcal{L}(Y)$-a.e. In this case, we will typically omit the $\sigma$, and write, for instance, that $\mathcal{L}(X\mid Y)=\mu(Y)$. Note that for fixed $y\in T$, the probability measure $\mu(y,\cdot)$ is not uniquely determined, since the kernel $\mu$ is only unique $\mathcal{L}(Y)$-a.e. Nonetheless, if a particular $\mu$ has been fixed, we will use the notation $\mathcal{L}(X\mid Y=y)$ to denote the probability measure $\mu(y,\cdot)$.
2.3 Exchangeability
A sequence $\xi=\{\xi_{i}:i\in\mathbb{N}\}$ of $S$-valued random variables is conditionally i.i.d. if there exists a $\sigma$-algebra $\mathcal{G}\subseteq\mathcal{F}$ and a random probability measure $\nu$ on $S$ such that $P(\xi\in A\mid\mathcal{G})=\nu^{\infty}(A)$ a.s., for all $A\in\mathcal{S}^{\infty}$. It can be shown that a sequence $\xi$ is conditionally i.i.d. if and only if there is a random probability measure $\mu$ such that
[…]
and in this case, $\mu$ can be chosen so that $\mu\in\mathcal{G}$ and $\mu=\nu$ a.s. (see [4, Lemma 2.1]).
A sequence $\xi=\{\xi_{i}:i\in\mathbb{N}\}$ of $S$-valued random variables is exchangeable if
[…]
whenever $k_{1},\ldots,k_{m}$ are distinct elements of $\mathbb{N}$. A function $\sigma:\mathbb{N}\to\mathbb{N}$ is called a (finite) permuatation if $\sigma$ is bijective and there exists $n_{0}\in\mathbb{N}$ such that $\sigma(n)=n$ for all $n\geq n_{0}$. The sequence $\xi$ is exchangeable if and only if $\{\xi_{\sigma(i)}\}$ and $\{\xi_{i}\}$ have the same finite-dimensional distributions whenever $\sigma$ is a permutation.
By [8, Theorem 1.1], we have that $\xi$ is conditionally i.i.d. if and only if $\xi$ is exchangeable. In that case, the random measure $\mu$ in (2.1) is a.s. unique, $\sigma(\xi)$-measurable, and satisfies $\mu_{n}(B)\to\mu(B)$ a.s., for all $B\in\mathcal{S}$, where
$\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\xi_{i}}$
are the empirical measures (see [8, Proposition 1.4]). In fact, we have that $\mu_{n}\to\mu$ a.s. in $M_{1}$ under the Prohorov metric (see [4, Proposition 2.4]). In [4, Proposition 2.5], the following conditional version of the law of large numbers is presented.
Proposition 2.1.
Let $(T,\mathcal{T})$ be a measurable space. Suppose $Y$ is a $T$-valued random variable, $\xi$ is conditionally i.i.d. given $\mathcal{G}$, and $f:T\times S\to\mathbb{R}$ is a measurable function such that $E|f(Y,\xi_{1})|<\infty$. If $Y\in\mathcal{G}$ and $Z$ is any version of $E[f(Y,\xi_{1})\mid\mathcal{G}]$ then
[…]
In particular,
[…]
2.4 Bayesian inference for row exchangeable arrays
The following four theorems come from [4]. The first is a de Finetti theorem for row exchangeability ([4, Theorem 3.2]). The final three ([4, Theorems 3.3, 3.4, 3.5]) provide needed tools for performing Bayesian inference on row exchangeable arrays.
Theorem 2.2.
Suppose $\xi$ is a row exchangeable, infinite array of $S$-valued random variables. Then there exists an a.s. unique random probability measure $\varpi$ on $M_{1}$ and an a.s. unique sequence $\mu=\{\mu_{i}:i\in\mathbb{N}\}$ of random probability measures on $S$ such that ${\mu\mid\varpi}\sim\varpi^{\infty}$ and ${\xi_{i}\mid\mu_{i}}\sim\mu_{i}^{\infty}$ for all $i\in\mathbb{N}$.
Theorem 2.3.
If $\xi$ is row exchangeable, then
[…]
for all $m,n\in\mathbb{N}$.
Theorem 2.4.
Let $\xi$ be row exchangeable. Fix $M\in\mathbb{N}$. For each $i\in\{1,\ldots,M\}$, let $N_{i},O_{i}\in\mathbb{N}$ with $N_{i}<O_{i}$, and let $A_{ij}\in\mathcal{S}$ for $i\leq M$ and $N_{i}<j\leq O_{i}$. Let $\mathcal{G}$ be a sub-$\sigma$-algebra of $\sigma(X_{1N_{1}},X_{2N_{2}},\ldots,X_{MN_{M}})$. Then
$P\left(\bigcap_{i=1}^{M}\bigcap_{j=N_{i}+1}^{O_{i}}\{\xi_{ij}\in A_{ij}\}\;\middle|\;\mathcal{G}\right)=E\left[\prod_{i=1}^{M}\prod_{j=N_{i}+1}^{O_{i}}\mu_{i}(A_{ij})\;\middle|\;\mathcal{G}\right].$
Theorem 2.5.
Let $\xi$ be row exchangeable. Fix $m,M,N\in\mathbb{N}$ with $m<M$. Then ${\bm{X}}_{mN}$ and $\{(\mu_{j},X_{jN})\}_{j=m+1}^{M}$ are conditionally independent given $\bm{\mu}_{m}$.
According to Theorem 2.4 with $N_{i}=N$, $O_{i}=N^{\prime}$, and $\mathcal{G}=\sigma({\bm{X}}_{MN})$, we can compute the posterior distribution (1.2), provided we can compute $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. Note that
$\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN};d\nu)=\mathcal{L}(\mu_{2},\ldots,\mu_{M}\mid{\bm{X}}_{MN},\mu_{1}=\nu_{1};d\nu_{2}\cdots d\nu_{M})\,\mathcal{L}(\mu_{1}\mid{\bm{X}}_{MN};d\nu_{1}).$
Iterating this, we see that we can determine $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$ if we know
$\mathcal{L}(\mu_{m}\mid{\bm{X}}_{MN},\bm{\mu}_{m-1}),\quad 1\leq m\leq M.$ (2.3)
Theorem 2.5 shows that in (2.3), the first $m-1$ rows of ${\bm{X}}_{MN}$ can be omitted. Hence, the conditional distribution (1.2) is entirely determined by the conditional distributions given in (1.3).
3 Overview of the Dirichlet process
3.1 Definition
The gamma distribution on $\mathbb{R}$ is defined by
$\operatorname{Gamma}(\alpha,\lambda)(dx)\propto 1_{(0,\infty)}x^{\alpha-1}e^{\lambda x}\,dx$
for $\alpha>0$ and $\lambda>0$, and $\operatorname{Gamma}(0,\lambda)=\delta_{0}$ for $\lambda>0$. The beta distribution on $\mathbb{R}$ is defined by
$\operatorname{Beta}(\alpha,\beta)(dx)\propto 1_{(0,1)}x^{\alpha-1}(1-x)^{\beta-1}\,dx$
for $\alpha>0$ and $\beta>0$, $\operatorname{Beta}(0,\beta)=\delta_{0}$ for $\beta>0$, and $\operatorname{Beta}(\alpha,0)=\delta_{1}$ for $\alpha>0$.
Let $\alpha=(\alpha_{0},\ldots,\alpha_{d})\in[0,\infty)^{d+1}$ with $\alpha_{j}>0$ for some $j$. Let $Z_{0},\ldots,Z_{d}$ be independent with $Z_{j}\sim\operatorname{Gamma}(\alpha_{j},1)$, and define $Y_{j}=Z_{j}/(\sum_{i=1}^{k}Z_{i})$. Then the Dirichlet distribution with parameter $\alpha$ is defined to be the distribution of $(Y_{0},\ldots,Y_{d})$. We denote it by $\operatorname{Dir}(\alpha)$. It is a probability measure on the simplex, $\Delta^{d}$, defined in (1.1).
If $\alpha_{j}=0$, then $Y_{j}=0$ a.s. If $\alpha_{j}>0$ for all $j$, then $\operatorname{Dir}(\alpha)$ is the probability measure on $\Delta^{d}$ proportional to $t_{0}^{\alpha_{0}-1}\cdots t_{d}^{\alpha_{d}-1}\,dt$. Here, $dt$ denotes the surface measure on $\Delta^{d}$. For fixed $j$, we have $Y_{j}\sim\operatorname{Beta}(\alpha_{j},\sum_{i\neq j}\alpha_{i})$.
Given a nonzero, finite measure $\alpha$ on $S$, we define a Dirichlet process on $S$ with parameter $\alpha$ to be a random probability measure $\lambda$ on $S$ that satisfies
$\mathcal{L}(\lambda(B_{0}),\ldots,\lambda(B_{d}))=\operatorname{Dir}(\alpha(B_{0}),\ldots,\alpha(B_{d})),$ (3.1)
whenever $\{B_{0},\ldots,B_{d}\}\subseteq\mathcal{S}$ is a partition of $S$. Let $\mathcal{D}(\alpha)$ denote the law of a Dirichlet process with parameter $\alpha$. Since a Dirichlet process is an $M_{1}$-valued random variable, it follows that $\mathcal{D}(\alpha)$ is a probability measure on $(M_{1},\mathcal{M}_{1})$. That is, $\mathcal{D}(\alpha)\in M_{1}(M_{1})$. If $B\in\mathcal{M}_{1}$, then we write $\mathcal{D}(\alpha,B)$ for $(\mathcal{D}(\alpha))(B)$.
Given $\alpha$ as above, let $\kappa=\alpha(S)>0$ and $\rho=\kappa^{-1}\alpha$, so that $\rho\in M_{1}$. We will typically write $\mathcal{D}(\alpha)=\mathcal{D}(\kappa\rho)$, and think of the law of a Dirichlet process as being determined by two parameters, a positive number $\kappa\in(0,\infty)$ and a probability measure $\rho\in M_{1}$. The measure $\rho$ is called the base measure, or base distribution, and the number $\kappa$ is called the concentration parameter. The following theorem shows that $\rho$ is the mean measure of $\lambda$ and can be found in [5, Theorem 4.3].
Theorem 3.1.
Let $\varphi:S\to\mathbb{R}$ be measurable. If $\varphi\in L^{1}(\rho)$, then
[…]
Taking $\varphi=1_{A}$ gives the special case, $E[\lambda(A)]=\rho(A)$ for all $A\in\mathcal{S}$.
Written analytically, (3.2) takes the form
$\int_{M_{1}}\int_{S}\varphi(x)\,\nu(dx)\,\mathcal{D}(\kappa\rho,d\nu)=\int_{S}\varphi(x)\,\rho(dx).$
Taking $\varphi=1_{A}$, we have
$\int_{M_{1}}\nu(A)\,\mathcal{D}(\kappa\rho,d\nu)=\rho(A),$ (3.3)
for all $A\in\mathcal{S}$.
3.2 The stick-breaking construction
The Dirichlet process can be built explicitly through the Sethuraman stick-breaking construction, which we outline below. See [14] for further details.
Let $V=\{V_{k}\}_{k=1}^{\infty}$ be i.i.d. $(0,1)$-valued random variables with $V_{k}\sim\operatorname{Beta}(1,\kappa)$. (If $\kappa=1$, then each $V_{k}$ is uniformly distributed. The larger $\kappa$ is, the smaller the $V_{k}$’s will tend to be; and conversely, the smaller $\kappa$ is, the larger the $V_{k}$’s will tend to be.) Note that by the second Borel-Cantelli lemma, $\sum_{k}V_{k}=\infty$ a.s. Choose $A\in\mathcal{F}$ such that $P(A)=1$ and $\sum_{k}V_{k}(\omega)=\infty$ for all $\omega\in A$.
We define a sequence $R=\{R_{k}\}_{k=1}^{\infty}$ of random variables as follows. On $A^{c}$, let $R_{1}=1$ and $R_{k}=0$ for $k>1$. On $A$, recursively define $R_{1}=V_{1}$ and
$R_{k+1}=\left(1-\sum_{j=1}^{k}R_{j}\right)V_{k+1}.$
The $R_{k}$’s can be thought of as the lengths of pieces of a stick of length 1. Initially, we break off a piece of length $V_{1}$. After we have broken off $n$ pieces, we take what remains and break off a proportion of it, where $V_{k+1}$ is that proportion. On $A$, the $R_{k}$’s can be expressed explicitly as
$R_{k}=V_{k}\prod_{j=1}^{k-1}(1-V_{j}).$
Note that $\sum_{k=1}^{\infty}R_{k}=1$. This is trivially the case on $A^{c}$. To see that it also holds on $A$, note that the above two displays give
$1-\sum_{j=1}^{k}R_{j}=\frac{R_{k+1}}{V_{k+1}}=\prod_{j=1}^{k}(1-V_{j}).$
Using results about infinite products, this shows that $0<\sum_{j=1}^{\infty}R_{j}\leq 1$, and $\sum_{j=1}^{\infty}R_{j}=1$ if and only if $\sum_{j=1}^{\infty}V_{j}=\infty$.
Let $\{U_{k}\}_{k=1}^{\infty}$ be i.i.d. $S$-valued random variables with $U_{k}\sim\rho$, and define
$\lambda=\sum_{k=1}^{\infty}R_{k}\delta_{U_{k}},$
so that $\lambda:\Omega\to M_{1}$ is a random probability measure that satisfies (3.1).
Note that if $\kappa$ is small, then the maximum of the $R_{k}$’s will tend to be large, meaning that $\lambda$ is likely to be concentrated on only a few points in $S$. Conversely, if $\kappa$ is large, then the $R_{k}$’s will be small, so that the mass of $\lambda$ is spread out over many points in $S$.
It can be shown, using the stick-breaking construction, that the function $(\kappa,\rho)\mapsto\mathcal{D}(\kappa\rho)$ is continuous and therefore a kernel from $(0,\infty)\times M_{1}$ to $M_{1}$. Moreover, if $\rho$ is not a point mass, then $\mathcal{D}(\kappa\rho)$ has no discrete component. That is, $\mathcal{D}(\kappa\rho,\{\nu\})=0$ for all $\nu\in M_{1}$.
3.3 Samples from a Dirichlet process
A sequence of samples from a Dirichlet process $\lambda\sim\mathcal{D}(\kappa\rho)$ is a sequence $\eta=\{\eta_{i}\}_{i=1}^{\infty}$ that satisfies ${\eta\mid\lambda}\sim\lambda^{\infty}$. We adopt the notation $\bm{\eta}_{n}=(\eta_{1},\ldots,\eta_{n})$ and ${\bm{x}}_{n}=(x_{1},\ldots,x_{n})\in S^{n}$. Note that for fixed $i$, we have
$P(\eta_{i}\in A)=E[P(\eta_{i}\in A\mid\lambda)]=E[\lambda(A)]=\rho(A).$ (3.4)
Thus, $\rho$ represents our prior distribution on the individual $\eta_{i}$’s, in the case that we have not observed any of their values. The posterior distribution is given in the following theorem, which can be found in [5, Theorem 3.1].
Theorem 3.2.
With notation as above, we have
$\mathcal{L}(\lambda\mid\bm{\eta}_{n})$ $=\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}}\bigg{)},\text{ and}$ (3.5)
$\mathcal{L}(\eta_{n+1}\mid\bm{\eta}_{n})$ $=\frac{\kappa}{\kappa+n}\rho+\frac{n}{\kappa+n}\widehat{\rho}_{n},$ (3.6)
where $\widehat{\rho}_{n}=n^{-1}\sum_{i=1}^{n}\delta_{\eta_{i}}$ is the empirical distribution of $\bm{\eta}_{n}$.
The next proposition expresses (3.5) in a purely analytic form.
Proposition 3.3.
Let $\alpha$ be a nonzero, finite measure on $S$. Then
$\int_{B}\nu^{n}(A)\,\mathcal{D}(\alpha,d\nu)=\int_{M_{1}}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\nu^{n}(d{\bm{x}}_{n})\,\mathcal{D}(\alpha,d\nu),$ (3.7)
for every $A\in\mathcal{S}^{n}$ and every $B\in\mathcal{M}_{1}$.
Proof.
Let $\lambda$ and $\eta$ be as in Theorem 3.2. Then
$P(\lambda\in B,\bm{\eta}_{n}\in A)=E[1_{B}(\lambda)P(\bm{\eta}_{n}\in A\mid\lambda)]=E[1_{B}(\lambda)\lambda^{n}(A)]=\int_{B}\nu^{n}(A)\,\mathcal{D}(\alpha,d\nu).$
On the other hand, by (3.5), we have
$P(\lambda\in B,\bm{\eta}_{n}\in A)$ $=E[1_{A}(\bm{\eta}_{n})P(\lambda\in B\mid\bm{\eta}_{n})]$
$=E\bigg{[}1_{A}(\bm{\eta}_{n})\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}\bigg{]}$
$=E\bigg{[}E\bigg{[}1_{A}(\bm{\eta}_{n})\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}\;\bigg{|}\;\lambda\bigg{]}\bigg{]}$
$=E\bigg{[}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\lambda^{n}(d{\bm{x}}_{n})\bigg{]}$
$=\int_{M_{1}}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\nu^{n}(d{\bm{x}}_{n})\,\mathcal{D}(\alpha,d\nu),$
which proves (3.7). ∎
We can also use (3.6) to obtain a recursive formula for the distribution of $\bm{\eta}_{n}$. Let $\rho_{n}=\mathcal{L}(\bm{\eta}_{n})$. Suppose $f:S^{n+1}\to\mathbb{R}$ is bounded and measurable. Then
[…]
Using (3.6), this gives
[…]
Hence,
[…]
3.4 Mixtures and noisy observations
[…]
Let $\alpha$ be a nonzero, finite random measure on $S$. That is, $\alpha$ is a random measure on $S$ such that $\alpha(S)\in(0,\infty)$ a.s. Let $\lambda$ be a random probability measure on $S$ such that $\lambda\mid\alpha\sim\mathcal{D}(\alpha)$. (This notation is meaningful, since $\alpha\mapsto\mathcal{D}(\alpha)$ is a probability kernel.) In this case we call $\lambda$ a mixture of Dirichlet processes on $S$ with mixing distribution $\mathcal{L}(\alpha)$. We also let $\kappa=\alpha(S)$ and $\rho=\alpha/\alpha(S)$, so that $\kappa$ and $\rho$ are random variables taking values in $(0,\infty)$ and $M_{1}$, respectively.
Let $\lambda$ be a mixture of Dirichlet processes as above. A sequence of samples from $\lambda$ is a sequence $\eta=\{\eta_{i}\}_{i=1}^{\infty}$ that satisfies ${\eta\mid\lambda,\alpha}\sim\lambda^{\infty}$. The following proposition generalizes (3.5) to mixtures.
Proposition 3.4.
With notation as above, we have
$\mathcal{L}(\lambda\mid\bm{\eta}_{n},\alpha)=\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}}\bigg{)},$
for any $n\in\mathbb{N}$.
Proof.
Let $A\in\mathcal{S}^{n}$ and $B,C\in\mathcal{M}_{1}$. Using Proposition 3.3, we have
$P(\lambda\in B,\bm{\eta}_{n}\in A,\alpha\in C)$ $=E[1_{C}(\alpha)1_{B}(\lambda)P(\bm{\eta}_{n}\in A\mid\lambda,\alpha)]$
$=E[1_{C}(\alpha)1_{B}(\lambda)\lambda^{n}(A)]$
$=E[1_{C}(\alpha)E[1_{B}(\lambda)\lambda^{n}(A)\mid\alpha]]$
$=E\bigg{[}1_{C}(\alpha)\int_{B}\nu^{n}(A)\,\mathcal{D}(\alpha,d\nu)\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)\int_{M_{1}}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\nu^{n}(d{\bm{x}}_{n})\,\mathcal{D}(\alpha,d\nu)\bigg{]}.$
On the other hand,
$E\bigg{[}\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}1_{A}(\bm{\eta}_{n})\,$ $1_{C}(\alpha)\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)E\bigg{[}1_{A}(\bm{\eta}_{n})\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}\;\bigg{|}\;\lambda,\alpha\bigg{]}\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\lambda^{n}(d{\bm{x}}_{n})\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)E\bigg{[}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\lambda^{n}(d{\bm{x}}_{n})\;\bigg{|}\;\alpha\bigg{]}\bigg{]}$
$=E\bigg{[}1_{C}(\alpha)\int_{M_{1}}\int_{A}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\nu^{n}(d{\bm{x}}_{n})\,\mathcal{D}(\alpha,d\nu)\bigg{]}.$
Hence,
[…]
which implies $P(\lambda\in B\mid\bm{\eta}_{n},\alpha)=\mathcal{D}(\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B)$. ∎
Now let $(T,\mathcal{T})$ be a measurable space. Fix $n\in\mathbb{N}$ and let $Y$ be a $T$-valued random variable such that $Y$ and $(\lambda,\alpha)$ are conditionally independent given $\bm{\eta}_{n}$. This holds, for example, if $Y$ is a function of $\bm{\eta}_{n}$ and $W$, where $W$ is some noise that is independent of $(\lambda,\alpha)$. In other words, we can think of $Y$ as a noisy observation of $\bm{\eta}_{n}$.
The following result extends (3.5) to noisy observations of data generated by a Dirichlet mixture. A special case of this appears as [1, Theorem 3].
Theorem 3.5.
With notation as above, we have
$\mathcal{L}(\lambda\mid Y)=\int_{(0,\infty)\times M_{1}\times S^{n}}{\mathcal{D}\bigg{(}t\nu+\sum_{i=1}^{n}\delta_{x_{i}}\bigg{)}}\,\mathcal{L}(\kappa,\rho,\bm{\eta}_{n}\mid Y;dt\,d\nu\,dx),$ (3.9)
In particular, if $\alpha$ is not random, as in the setting of Theorem 3.2, so that $Y$ and $\lambda$ are conditionally independent given $\bm{\eta}_{n}$, then
$\mathcal{L}(\lambda\mid Y)=\int_{S^{n}}{\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{x_{i}}\bigg{)}}\,\mathcal{L}(\bm{\eta}_{n}\mid Y;d{\bm{x}}_{n})$ (3.10)
Proof.
Let $B\in\mathcal{M}_{1}$. Then
$P(\lambda\in B\mid Y)=E[P(\lambda\in B\mid Y,\bm{\eta}_{n})\mid Y]=E[P(\lambda\in B\mid\bm{\eta}_{n})\mid Y],$ (3.11)
since $Y$ and $\lambda$ are independent given $\bm{\eta}_{n}$. Similarly,
$P(\lambda\in B\mid\bm{\eta}_{n})=E[P(\lambda\in B\mid\bm{\eta}_{n},\alpha)\mid\bm{\eta}_{n}]=E[P(\lambda\in B\mid\bm{\eta}_{n},\alpha)\mid Y,\bm{\eta}_{n}],$ (3.12)
since $Y$ and $\alpha$ are independent given $\bm{\eta}_{n}$. Substituting (3.12) in (3.11), we have
$P(\lambda\in B\mid Y)=E[E[P(\lambda\in B\mid\bm{\eta}_{n},\alpha)\mid Y,\bm{\eta}_{n}]\mid Y]=E[P(\lambda\in B\mid\bm{\eta}_{n},\alpha)\mid Y].$
By Proposition 3.4, this gives
$P(\lambda\in B\mid Y)$ $=E\bigg{[}\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}},B\bigg{)}\;\bigg{|}\;Y\bigg{]}$
$=\int_{(0,\infty)\times M_{1}\times S^{n}}{\mathcal{D}\bigg{(}t\nu+\sum_{i=1}^{n}\delta_{x_{i}},B\bigg{)}}\,\mathcal{L}(\kappa,\rho,\bm{\eta}_{n}\mid Y;dt\,d\nu\,dx),$
which is (3.9). In the case that $\alpha$ is not random, this reduces to (3.10). ∎
3.5 Understanding the IDP parameters
Let $\varpi\sim\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$ be an IDP and $\xi=\{\xi_{mn}\}_{m,n\in\mathbb{N}}$ an array of samples from $\varpi$. By this we mean there exists a sequence of random measures $\mu=\{\mu_{m}\}_{m\in\mathbb{N}}$ such that ${\mu\mid\varpi}\sim\varpi^{\infty}$ and ${\xi_{m}\mid\mu_{m}}\sim\mu_{m}^{\infty}$. Note that each $\mu_{m}$ is a function of $\xi_{m}$. More specifically, the random measure $\mu_{m}$ is the almost sure weak limit of the empirical distributions $n^{-1}\sum_{j=1}^{n}\delta_{\xi_{mj}}$. As noted in Section 1, we call the $\mu_{m}$’s the row distributions, and we call $\varpi$ the row distribution generator.
Since ${\xi_{m}\mid\mu_{m}}\sim\mu_{m}^{\infty}$ and $\mu_{m}\sim\mathcal{D}(\varepsilon\varrho)$, we may apply the results of Section 3.3. For example, by (3.4), we have $P(\xi_{mn}\in A)=\varrho(A)$. In other words, the base measure $\varrho$ is the prior distribution on the samples $\xi_{mn}$.
Similarly, by (3.6), we have
$P(\xi_{m,n+1}\in A\mid X_{mn})=\frac{\varepsilon}{\varepsilon+n}\varrho(A)+\frac{1}{\varepsilon+n}\sum_{j=1}^{n}\delta_{\xi_{mj}}(A).$
As in Section 3.3, if the row concentration $\varepsilon$ is small, then our prior expectation on the empirical distributions of the rows $\xi_{m}$ is that they are concentrated on only a few points in $S$. Conversely, if $\varepsilon$ is large, then we expect those empirical distributions to be spread out over many points.
By Theorem 2.3 and (3.6) we have
$P(\xi_{m+1,n}\in A\mid\bm{\mu}_{m})$ $=E[P(\xi_{m+1,n}\in A\mid\bm{\mu}_{m+1})\mid\bm{\mu}_{m}]$
$=E[\mu_{m+1}(A)\mid\bm{\mu}_{m}]$
$=\int_{M_{1}}\nu(A)\,\mathcal{L}(\mu_{m+1}\mid\bm{\mu}_{m};d\nu)$
$=\frac{\kappa}{\kappa+m}\int_{M_{1}}\nu(A)\,\mathcal{D}(\varepsilon\varrho,d\nu)+\frac{1}{\kappa+m}\sum_{i=1}^{m}\mu_{i}(A).$
Hence, (3.3) gives
$P(\xi_{m+1,n}\in A\mid\bm{\mu}_{m})=\frac{\kappa}{\kappa+m}\varrho(A)+\frac{1}{\kappa+m}\sum_{i=1}^{m}\mu_{i}(A).$
As above, if the column concentration $\kappa$ is small, then we expect the empirical distribution of the $\mu_{m}$’s to be concentrated on only a few points in $M_{1}$. In other words, we expect the $\mu$’s to take on only a handful of different possible values. Conversely, if $\kappa$ is large, then we expect the $\mu_{m}$’s to take on many different values.
Remark 3.6.
The IDP bears some superficial resemblance to what is called a hierarchical Dirichlet process. A hierarchical Dirichlet process is a mixture of Dirichlet processes on $S$ with mixing distribution $\mathcal{L}(\alpha)$, where $\alpha/\alpha(S)$ is itself a Dirichlet process (see [15, 16]).
An IDP is a Dirichlet process on $M_{1}$ whose base measure is the law of a Dirichlet process. A hierarchical Dirichlet process, on the other hand, is as a Dirichlet process on $S$ whose base measure is a Dirichlet process itself. These are actually quite different. In fact, the two processes have different state spaces. A hierarchical Dirichlet process takes values in $M_{1}$, whereas an IDP takes values in $M_{1}(M_{1})$.
For example, if we take $\kappa^{\prime}$ to be random and $\rho^{\prime}\sim\mathcal{D}(\varepsilon\varrho)$, then we can define a hierarchical Dirichlet process $\lambda$ by the conditional distribution ${\lambda\mid\kappa^{\prime},\rho^{\prime}}\sim\mathcal{D}(\kappa^{\prime}\rho^{\prime})$. Note that $\rho^{\prime}$ is an $M_{1}$-valued random variable.
In contrast, for the IDP, we take $\kappa$ and $\rho$ to be nonrandom and $\rho=\mathcal{D}(\varepsilon\varrho)$. In this case, $\rho$ is a nonrandom element of $M_{1}(M_{1})$. We then define an IDP $\varpi$ by the unconditional distribution $\varpi\sim\mathcal{D}(\kappa\rho)$.
4 Sequential imputation
Recall that our goal is to compute (1.2), which we can do by computing (1.3). As we saw in Section 1, it will be unreasonable to calculate (1.3) exactly, at least when $m<M$, since in that case we would need to simultaneously condition on multiple rows of observations. If, however, we condition on only one row of observations, then we can make progress. (See Section 5.3.) Suppose, for instance, that we can compute (1.4) for $1\leq m\leq M$. If we can use these instead of (1.3) to compute (1.2), then we can avoid the issues that arose in the exact calculation of (1.3).
By Theorem 2.4, we must find a way to use (1.4) to compute $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. If we do this via simulation, then we must find a way to use (1.4) to simulate $\bm{\mu}_{M}$ according to the conditional distribution $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$.
One approach would be to simulate $\mu_{1}$ according to the distribution $\mathcal{L}(\mu_{1}\mid X_{1N})$, and then use that simulated value to simulate $\mu_{2}$ according to $\mathcal{L}(\mu_{2}\mid X_{2N},\mu_{1})$, and so on. However, if we do that, then we would not be simulating $\bm{\mu}_{M}$ according to its correct conditional distribution, since our simulation of $\mu_{m}$ would not take into account observations from higher-numbered rows.
One way to fix this is to do many such incorrect simulations of $\bm{\mu}_{M}$. Let $K$ be the number of incorrect simulations we generate. Some of these $K$ simulations will be “more incorrect” than others. We then assign the $K$ simulated values weights according to their level of correctness, and choose one of them randomly, with probabilities proportional to those weights. If we assign the weights appropriately, then the distribution of the chosen value will converge to $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$ as $K$ tends to infinity.
This is the method of sequential imputation, first introduced in [11]. It is an application of the more general method of importance sampling that originated in [9]. In this section, we show how to apply sequential imputation to the IDP. This was done in [12] in the special case $S=\{0,1\}$. The work in [12] was based on the work in [11], which evidently assumed that the underlying random variables have densities with respect to some product measure. In Proposition 4.5, we show that this assumption is not satisfied by the IDP. In particular, the work in [11] is not sufficient to justify its use in [12]. In Theorem 4.7, we provide a new proof that justifies this approach in general, including the case $S=\{0,1\}$ treated in [12] as well as the general case treated here.
4.1 Importance sampling
Importance sampling is a method of approximating a particular probability distribution using samples from a different distribution. The samples themselves will vary in how “important” they are in determining the distribution of interest. This is modeled by assigning different weights to the samples.
We begin by presenting, without commentary, the formal statement of the method of importance sampling in Theorem 4.1 below. We then describe in Remark 4.2 the intuitive interpretation of the method.
Let $(T,\mathcal{T})$ be a measurable space. Let $Z$ be an $S$-valued random variable and $Y$ a $T$-valued random variable. Let $\mathfrak{m}^{*}$ be a probability kernel from $T$ to $S$ such that $\mathcal{L}(Z\mid Y)\ll\mathfrak{m}^{*}(Y)$ a.s. Assume there exist measurable functions $w:T\times S\to\mathbb{R}$ and $h:T\to[0,\infty)$ such that $h(Y)>0$ a.s., $Eh(Y)<\infty$, and
[…]
Define $Z^{*}$ so that $Z^{*}\mid Y\sim\mathfrak{m}^{*}(Y)$ and let $W=w(Y,Z^{*})$. Let $\{(Z^{*,k},W_{k})\}_{k=1}^{\infty}$ be copies of $(Z^{*},W)$ that are i.i.d. given $Y$. Define $\widetilde{Z}^{K}$ so that
$\widetilde{Z}^{K}\mid Z^{*,1},\ldots,Z^{*,K},Y\propto\sum_{k=1}^{K}{W_{k}\delta_{Z^{*,k}}}$ (4.2)
Theorem 4.1.
With the notation and assumptions given above, we have
[…]
as $K\to\infty$.
Remark 4.2.
The interpretation of Theorem 4.1 is the following. We observe $Y$ and we wish to determine $\mathcal{L}(Z\mid Y)$. Unfortunately, for one reason or another, this is not directly possible. Instead, we are only able to determine a different distribution, $\mathfrak{m}^{*}(Y)$, which we call the simulation measure. Using $\mathfrak{m}^{*}(Y)$, we generate an i.i.d. collection of samples, $Z^{*,1},\ldots,Z^{*,K}$. The $k$-th sample, $Z^{*,k}$, gets assigned the weight $W_{k}=w(Y,Z^{*,k})$, where $w$ is some function satisfying (4.1). We then use these weights to randomly choose one of the $K$ samples. The randomly chosen sample is denoted by $\widetilde{Z}^{K}$. Theorem 4.1 says that if $K$ is large, then the law of $\widetilde{Z}^{K}$ is close to $\mathcal{L}(Z\mid Y)$.
Proof of Theorem 4.1.
First note that
$E[w(Y,Z^{*,1})\mid Y]=\int_{S}w(Y,z)\,\mathfrak{m}^{*}(Y,dz)=h(Y).$ (4.4)
Hence, $Ew(Y,Z^{*,1})=Eh(Y)<\infty$. Now let $f:S\to\mathbb{R}$ be bounded and measurable. By Proposition 2.1, we have
[…]
But
$E[w(Y,Z^{*,1})f(Z^{*,1})\mid Y]=\int_{S}w(Y,z)f(z)\,\mathfrak{m}^{*}(Y,dz)\\ =h(Y)\int_{S}f(z)\,\mathcal{L}(Z\mid Y;dz)=h(Y)E[f(Z)\mid Y].$
Therefore, since $h(Y)>0$ a.s.,
$E[f(\widetilde{Z}^{K})\mid Z^{*,1},\ldots,Z^{*,K},Y]$ $=\frac{\sum_{k=1}^{K}w(Y,Z^{*,k})f(Z^{*,k})}{\sum_{k=1}^{K}w(Y,Z^{*,k})}$
$\to\frac{h(Y)E[f(Z)\mid Y]}{h(Y)E[1\mid Y]}$
$=E[f(Z)\mid Y],$
and this proves (4.3). ∎
4.2 Effective sample size
Let $f:S\to\mathbb{R}$ be continuous and bounded. By (4.3), if $K$ is large, then
$\frac{\sum_{k=1}^{K}W_{k}f(Z^{*,k})}{\sum_{k=1}^{K}W_{k}}\approx E[f(Z)\mid Y].$
On the other hand, by (2.2), if $\{Z^{k}\}_{k=1}^{\infty}$ are copies of $Z$ that are i.i.d. given $Y$, then
$\frac{1}{K}\sum_{k=1}^{K}f(Z^{k})\approx E[f(Z)\mid Y].$
This latter estimate of $E[f(Z)\mid Y]$ is presumably more efficient, in the sense that smaller $K$ values are needed. This is because in the latter estimate, we are generating values directly from $\mathcal{L}(Z\mid Y)$, rather than from the modified distribution $\mathfrak{m}^{*}(Y)$.
In an effort to measure this difference in efficiency, let $K$ be a given number of weighted samples. We wish to find a number $K_{e}$ such that
$\operatorname{Var}\bigg{(}\frac{\sum_{k=1}^{K}W_{k}f(Z^{*,k})}{\sum_{k=1}^{K}W_{k}}\;\bigg{|}\;Y\bigg{)}\approx\operatorname{Var}\bigg{(}\frac{1}{K_{e}}\sum_{k=1}^{K_{e}}f(Z^{k})\;\bigg{|}\;Y\bigg{)}.$
The right-hand side is $K_{e}^{-1}\operatorname{Var}(f(Z)\mid Y)$. In [10], it is shown that
$\operatorname{Var}\bigg{(}\frac{\sum_{k=1}^{K}W_{k}f(Z^{*,k})}{\sum_{k=1}^{K}W_{k}}\;\bigg{|}\;Y\bigg{)}\approx\frac{1}{K}\operatorname{Var}(f(Z)\mid Y)\bigg{(}1+\operatorname{Var}\bigg{(}\frac{W}{h(Y)}\;\bigg{|}\;Y\bigg{)}\bigg{)}.$
We therefore define
$K_{e}=\frac{K}{1+\operatorname{Var}\left(\frac{W}{h(Y)}\;\middle|\;Y\right)},$
and call this the effective sample size.
By (4.4), we have
$\operatorname{Var}\left(\frac{W}{h(Y)}\;\middle|\;Y\right)=\frac{\operatorname{Var}(W\mid Y)}{h(Y)^{2}}=\frac{\operatorname{Var}(W\mid Y)}{E[W\mid Y]^{2}}.$
Therefore,
$K_{e}=K\left(1+\frac{\operatorname{Var}(W\mid Y)}{E[W\mid Y]^{2}}\right)^{-1}.$
If we approximate $E[W\mid Y]$ by the sample mean, $\overline{W}=K^{-1}\sum_{k=1}^{K}W_{k}$, and $\operatorname{Var}(W\mid Y)$ by the population variance, $\widetilde{S}^{2}=(K^{-1}\sum_{k=1}^{K}W_{k}^{2})-\overline{W}^{2}$, then we have $K_{e}\approx K_{e}^{\prime}$, where
$K_{e}^{\prime}=\frac{K}{1+\widetilde{S}^{2}/\overline{W}^{2}}=\frac{\big{(}\sum_{k=1}^{K}W_{k}\big{)}^{2}}{\sum_{k=1}^{K}W_{k}^{2}}$
On the other hand, if we use the sample variance, $S^{2}=K(K-1)^{-1}\widetilde{S}^{2}$, then we have $K_{e}\approx K_{e}^{\prime\prime}$, where
$K_{e}^{\prime\prime}=\frac{K}{1+S^{2}/\overline{W}^{2}}=\frac{K(K-1)}{K-1+K\widetilde{S}^{2}/\overline{W}^{2}}=\frac{K(K-1)}{K-1+K(K/K_{e}^{\prime}-1)}=\left(\frac{K-1}{K-K_{e}^{\prime}/K}\right)K_{e}^{\prime}.$
4.3 Sequential imputation and the simulation measure
Now fix $M\in\mathbb{N}$. Let $Z=(Z_{1},\ldots,Z_{M})$ be an $S^{M}$-valued random variable and let $z=(z_{1},\ldots,z_{M})$ denote an element of $S^{M}$. We adopt the notation ${\bm{Z}}_{m}=(Z_{1},\ldots,Z_{m})$ and we use ${\bm{z}}_{m}=(z_{1},\ldots,z_{m})$ for an element of $S^{m}$. Note that ${\bm{Z}}_{M}=Z$ and ${\bm{z}}_{M}=z$. We also let $Y=(Y_{1},\ldots,Y_{M})$ be a $T^{M}$-valued random variable and adopt similar notation in that case.
We think of $Y$ as observed values and $Z$ as unobserved. In this sense, $Z$ is regarded as “missing data.” We wish to determine $\mathcal{L}(Z\mid Y)$. Suppose, however, that we are only able to determine $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$ for $1\leq m\leq M$. (By convention, a variable with a 00 subscript is omitted. Hence, when $m=1$, we have $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})=\mathcal{L}(Z_{1}\mid Y_{1})$.) We describe here a method of using $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$ to approximate $\mathcal{L}(Z\mid Y)$. This method is called sequential imputation and first appeared in [11].
Consider, for the moment, the case $M=2$. By conditioning on $Z_{1}$, we could determine $\mathcal{L}(Z\mid Y)$ sequentially, if we could compute
  1. (i)
    $\mathcal{L}(Z_{1}\mid Y)$ and
  2. (ii)
    $\mathcal{L}(Z_{2}\mid Y,Z_{1})$.
The second of these is available to us, but the first is not. Instead of (i), we can only compute $\mathcal{L}(Z_{1}\mid Y_{1})$. The idea in sequential imputation is to use $\mathcal{L}(Z_{1}\mid Y_{1})$ to simulate $Z_{1}$, then use this simulated value in (ii) to determine the law of $Z_{2}$. We are substituting the missing data $Z_{1}$ with its (incorrectly) simulated value. This kind of substitution is called imputation. Since we are using $\mathcal{L}(Z_{1}\mid Y_{1})$ instead of the correct distribution in (i), we must combine this with the method of importance sampling presented in Theorem 4.1.
To apply Theorem 4.1, we first construct the simulation measure $\mathfrak{m}^{*}$. Let $\gamma_{m}$ be a probability kernel from $T^{m}\times S^{m-1}$ to $S$ with $Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1}\sim\gamma_{m}({\bm{Y}}_{m},{\bm{Z}}_{m-1})$. Let $\gamma_{m}^{*}$ be the probability kernel from $T^{M}\times S^{m-1}$ to $S$ given by $\gamma_{m}^{*}(y,{\bm{z}}_{m-1})=\gamma_{m}({\bm{y}}_{m},{\bm{z}}_{m-1})$. Note that $\gamma_{M-1}^{*}$ is a probability kernel from $T^{M}\times S^{M-2}$ to $S$ and $\gamma_{M}^{*}$ is a probability kernel from $T^{M}\times S^{M-1}$ to $S$. Hence, $\gamma_{M-1}^{*}\gamma_{M}^{*}$ is a probability kernel from $T^{M}\times S^{M-2}$ to $S^{2}$. Iterating this, if we define $\mathfrak{m}^{*}=\gamma_{1}^{*}\cdots\gamma_{M}^{*}$, then $\mathfrak{m}^{*}$ is a probability kernel from $T^{M}$ to $S^{M}$.
In Theorem 4.1, we have $Z^{*,k}\mid Y\sim\mathfrak{m}^{*}(Y)$. Hence,
$Z_{m}^{*,k}\mid Y,Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k}\sim\gamma_{m}^{*}(Y,Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k})=\gamma_{m}({\bm{Y}}_{m},Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k}).$ (4.5)
In other words, the simulated vector $Z^{*,k}=(Z_{1}^{*,k},\ldots,Z_{M}^{*,k})$ can be constructed sequentially using $\mathcal{L}(Z_{m}\mid{\bm{Y}}_{m},{\bm{Z}}_{m-1})$, where in each step the missing data ${\bm{Z}}_{m-1}$ is imputed with the previously simulated values $Z_{1}^{*,k},\ldots,Z_{m-1}^{*,k}$. To prove that Theorem 4.1 applies in this situation, we must find a weight function $w$ satisfying (4.1).
4.4 A simulation density
As noted earlier, sequential imputation first appeared in [11]. There, a weight function was constructed using density functions. The proof and construction in [11] did not specify the codomain of the random variables $Y$ and $Z$, nor did it specify the measures with respect to which they have joint and conditional densities. In Theorem 4.4 below, we give a rigorous formulation of the proof in [11]. First, we clarify the assumptions about the existence of densities, and then show how this relates to the simulation measure $\mathfrak{m}^{*}$.
Assumption 4.3.
There exist $\sigma$-finite measures $\mathfrak{n}$ and $\widetilde{\mathfrak{n}}$ on $S$ and $T$, respectively, such that $\mathcal{L}(Y,Z)\ll\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}^{M}$.
If Assumption 4.3 holds, then we may let $f=d\mathcal{L}(Y,Z)/d(\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}^{M})$ be a density of $(Y,Z)$ with respect to $\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}^{M}$. If we write $f$ with omitted arguments, it is assumed that they have been integrated out. For example,
$f(y_{1},{\bm{z}}_{m})=\int_{S^{M-m}}\int_{T^{M-1}}{f(y,z)}\,\widetilde{\mathfrak{n}}^{M-1}(dy_{2}\cdots dy_{M})\,\mathfrak{n}^{M-m}(dz_{m+1}\cdots dz_{m}).$
In other words, such functions are the marginal densities. By changing $f$ on a set of measure zero, we may assume $f\in[0,\infty)$ everywhere and if the value of a marginal density at a point is 00, then $f$ at that point is 00 for all values of the omitted arguments.
We use |||| to denote conditional densities. For example,
$f(z_{m+1}\mid{\bm{y}}_{m},{\bm{z}}_{m})=\frac{f({\bm{y}}_{m},{\bm{z}}_{m+1})}{f({\bm{y}}_{m},{\bm{z}}_{m})}.$
As usual, we adopt the convention that a variable with a 00 subscript is omitted. For instance, if $m=1$, then $f({\bm{y}}_{m},{\bm{z}}_{m-1})=f(y_{1})$ and $f(z_{m}\mid{\bm{y}}_{m},{\bm{z}}_{m-1})=f(z_{1}\mid y_{1})$.
With this notation, we may write
$\gamma_{m}^{*}(y,{\bm{z}}_{m-1},dz_{m})={f(z_{m}\mid{\bm{y}}_{m},{\bm{z}}_{m-1})\,\mathfrak{n}(dz_{m})}.$
We also have
$(\gamma_{M-1}^{*}\gamma_{M}^{*})(y,{\bm{z}}_{M-2},dz_{M-1}\,dz_{M})={\gamma_{M}^{*}(y,{\bm{z}}_{M-1},dz_{M})\,\gamma_{M-1}^{*}(y,{\bm{z}}_{M-2},dz_{M-1})}\\ ={f(z_{M}\mid{\bm{y}}_{M},{\bm{z}}_{M-1})f(z_{M-1}\mid{\bm{y}}_{M-1},{\bm{z}}_{M-2})\,\mathfrak{n}(dz_{M})\,\mathfrak{n}(dz_{M-1})}.$
Iterating this, we obtain $\mathfrak{m}^{*}(y,dz)=f^{*}(y,z)\,\mathfrak{n}^{M}(dz)$, where
$f^{*}(y,z)=\prod_{m=1}^{M}f(z_{m}\mid{\bm{y}}_{m},{\bm{z}}_{m-1}).$
4.5 A proof using densities
We now define the weight function and show that sequential imputation leads asymptotically to $\mathcal{L}(Z\mid Y)$. For $(y,z)\in T\times S$, define
$w(y,z)=\prod_{m=1}^{M}f(y_{m}\mid{\bm{y}}_{m-1},{\bm{z}}_{m-1}).$
Define $Z^{*,k}$ and $\widetilde{Z}^{K}$ as in (4.2).
Theorem 4.4.
If Assumption 4.3 holds and $f(y)\in L^{2}(\widetilde{\mathfrak{n}}^{M})$, then
[…]
Proof.
By Theorem 4.1, it suffices to show that $\mathcal{L}(Z\mid Y)\ll\mathfrak{m}^{*}(Y)$ a.s., $f(Y)>0$ a.s., $Ef(Y)<\infty$, and
[…]
Since $f(y)$ is the density of $Y$ with respect to $\widetilde{\mathfrak{n}}^{M}$, we have
$P(f(Y)=0)=\int_{f^{-1}(\{0\})}f(y)\,\widetilde{\mathfrak{n}}^{M}(dy)=0,$
so that $f(Y)>0$ a.s. Since $f(y)\in L^{2}(\widetilde{\mathfrak{n}}^{M})$, we also have
$Ef(Y)=\int_{T^{M}}f(y)^{2}\,\widetilde{\mathfrak{n}}^{M}(dy)<\infty.$
Finally,
$w(y,z)f^{*}(y,z)=\prod_{m=1}^{M}{\frac{f({\bm{y}}_{m},{\bm{z}}_{m-1})}{f({\bm{y}}_{m-1},{\bm{z}}_{m-1})}\,\frac{f({\bm{y}}_{m},{\bm{z}}_{m})}{f({\bm{y}}_{m},{\bm{z}}_{m-1})}}=f(y,z).$
Hence,
$\mathcal{L}(Z\mid Y=y;dz)=f(z\mid y)\,\mathfrak{n}^{M}(dz)=\frac{f(y,z)}{f(y)}\,\mathfrak{n}^{M}(dz)\\ =\frac{w(y,z)}{f(y)}\,f^{*}(y,z)\,\mathfrak{n}^{M}(dz)=\frac{w(y,z)}{f(y)}\,\mathfrak{m}^{*}(y,dz).$
Therefore, $d\mathcal{L}(Z\mid Y)/d\mathfrak{m}^{*}(Y)=w(Y,\cdot)/f(Y)$ a.s. ∎
4.6 The IDP has no density
We wish to apply sequential imputation to determine $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$, using the computable distributions (1.4). In this case, we would naturally take $Z_{m}=\mu_{m}$ and $Y_{m}=X_{mN}=(\xi_{m1},\xi_{m2},\ldots,\xi_{mN})$. In [12], the author did exactly this in the special case $S=\{0,1\}$.
Unfortunately, as we see below in Proposition 4.5, sequential imputation—as it is presented in Theorem 4.4—does not apply in this case. Namely, Assumption 4.3 is not satisfied. The vector $Z=\bm{\mu}_{m}$ has no joint density with respect to any product measure. Hence, the proof of Theorem 4.4, which is a rigorous presentation of the proof in [11], does not justify the use of sequential imputation in this setting. This includes not only the general setting that we are working with, but also the special case $S=\{0,1\}$ that was treated in [12].
Proposition 4.5.
If $\varrho$ is not a point mass, then there do not exist $\sigma$-finite measures $\mathfrak{n}_{1}$ and $\mathfrak{n}_{2}$ on $M_{1}$ such that $\mathcal{L}(\bm{\mu}_{2})\ll\mathfrak{n}_{1}\times\mathfrak{n}_{2}$.
Proof.
Suppose $\varrho$ is not a point mass. Assume that $\mathfrak{n}_{1}$ and $\mathfrak{n}_{2}$ are $\sigma$-finite measures on $M_{1}$ and $\mathcal{L}(\bm{\mu}_{2})\ll\mathfrak{n}_{1}\times\mathfrak{n}_{2}$. Let $A=\{\gamma\in M_{1}:\mathfrak{n}_{1}(\{\gamma\})>0\}$, so that $A$ is countable. Let $D=\{(\gamma,\gamma)\in M_{1}^{2}:\gamma\notin A\}$. Then
$(\mathfrak{n}_{1}\times\mathfrak{n}_{2})(D)=\int_{A^{c}}\mathfrak{n}_{1}(\{\gamma\})\,\mathfrak{n}_{2}(d\gamma)=0.$
Therefore, by hypothesis, $P(\bm{\mu}_{2}\in D)=0$. Since $\varrho$ is not a point mass, the measure $\rho=\mathcal{D}(\varepsilon\varrho)$ has no discrete component. Therefore, $P(\mu_{1}\in A)=\rho(A)=0$, since $A$ is countable. Hence, by [7, Theorem 5.4] and (3.6), we have
$P(\bm{\mu}_{2}\in D)=P(\mu_{1}\notin A,\mu_{1}=\mu_{2})=P(\mu_{1}=\mu_{2})=E[P(\mu_{1}=\mu_{2}\mid\mu_{1})]\\ =E[\mathcal{L}(\mu_{2}\mid\mu_{1};\{\mu_{1}\})]=E\bigg{[}\frac{\kappa}{\kappa+1}\rho(\{\mu_{1}\})+\frac{1}{\kappa+1}\delta_{\mu_{1}}(\{\mu_{1}\})\bigg{]}=\frac{1}{\kappa+1}>0,$
a contradiction. ∎
4.7 A proof without a simulation density
In order to prove that sequential imputation applies to the IDP, we must give a new proof based on a new assumption. We continue to let the simulation measure $\mathfrak{m}^{*}$ be defined as in Section 4.3. But we must drop Assumption 4.3, and consequently drop the assumption that $\mathfrak{m}^{*}$ has a density with respect to a product measure. We cannot drop densities altogether, though, since they are essential to defining the weight function.
Assumption 4.6.
There exist $\sigma$-finite measures $\mathfrak{n}_{1},\mathfrak{n}_{2},\ldots,\mathfrak{n}_{M}$ and $\widetilde{\mathfrak{n}}$ on $S,S^{2},\ldots,S^{M}$ and $T$, respectively, such that $\mathcal{L}(Y,{\bm{Z}}_{m})\ll\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}_{m}$ for every $m$.
Under Assumption 4.6, we may let $f_{m}$ be a density of $(Y,{\bm{Z}}_{m})$ with respect to $\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}_{m}$. We adopt the same assumptions and notational conventions for $f_{m}$ as we did for $f$ in Section 4.4. For $(y,z)\in T\times S$, define
$w(y,z)=f_{1}(y_{1})\prod_{m=1}^{M-1}{f_{m}(y_{m+1}\mid{\bm{y}}_{m},{\bm{z}}_{m})}.$ (4.6)
Define $Z^{*,k}$ and $\widetilde{Z}^{K}$ as in (4.2).
Theorem 4.7.
If Assumption 4.6 holds and $f_{M}(y)\in L^{2}(\widetilde{\mathfrak{n}}^{M})$, then
[…]
Proof.
The first part of the proof of Theorem 4.4 carries over, so we need only show that
[…]
Let $k\in\{1,\ldots,M-1\}$ and let $A\in\mathcal{T}$, $B\in\mathcal{S}^{k}$, and $C\in\mathcal{S}$. Then
$P(Y_{k+1}$ $\in A,{\bm{Z}}_{k}\in B,Z_{k+1}\in C\mid{\bm{Y}}_{k},{\bm{Z}}_{k})$
$=1_{B}({\bm{Z}}_{k})P(Y_{k+1}\in A,Z_{k+1}\in C\mid{\bm{Y}}_{k},{\bm{Z}}_{k})$
$=1_{B}({\bm{Z}}_{k})E[{P(Y_{k+1}\in A,Z_{k+1}\in C\mid{\bm{Y}}_{k+1},{\bm{Z}}_{k})}\mid{\bm{Y}}_{k},{\bm{Z}}_{k}]$
$=1_{B}({\bm{Z}}_{k})E[{1_{A}(Y_{k+1})\gamma_{k+1}({\bm{Y}}_{k+1},{\bm{Z}}_{k},C)}\mid{\bm{Y}}_{k},{\bm{Z}}_{k}]$
$=1_{B}({\bm{Z}}_{k})\int_{A}{\gamma_{k+1}({\bm{Y}}_{k},y_{k+1},{\bm{Z}}_{k},C)f_{k}(y_{k+1}\mid{\bm{Y}}_{k},{\bm{Z}}_{k})}\,\widetilde{\mathfrak{n}}(dy_{k+1}).$
Hence,
$P(Y_{k+1}$ $\in A,{\bm{Z}}_{k}\in B,Z_{k+1}\in C\mid{\bm{Y}}_{k})$
$=E\bigg{[}{1_{B}({\bm{Z}}_{k})\int_{A}{\gamma_{k+1}({\bm{Y}}_{k},y_{k+1},{\bm{Z}}_{k},C)f_{k}(y_{k+1}\mid{\bm{Y}}_{k},{\bm{Z}}_{k})}\,\widetilde{\mathfrak{n}}(dy_{k+1})}\bigg{|}{\bm{Y}}_{k}\bigg{]}$
$=\int_{B}\int_{A}{\gamma_{k+1}({\bm{Y}}_{k},y_{k+1},{\bm{z}}_{k},C)f_{k}(y_{k+1}\mid{\bm{Y}}_{k},{\bm{z}}_{k})}\,\widetilde{\mathfrak{n}}(dy_{k+1})f_{k}({\bm{z}}_{k}\mid{\bm{Y}}_{k})\,\mathfrak{n}_{k}(d{\bm{z}}_{k})$
$=\int_{A}\int_{B}\int_{C}{f_{k}(y_{k+1}\mid{\bm{Y}}_{k},{\bm{z}}_{k})\,\gamma_{k+1}({\bm{Y}}_{k},y_{k+1},{\bm{z}}_{k},dz_{k+1})f_{k}({\bm{z}}_{k}\mid{\bm{Y}}_{k})\,\mathfrak{n}_{k}(d{\bm{z}}_{k})\,\widetilde{\mathfrak{n}}(dy_{k+1})}.$
On the other hand,
$P(Y_{k+1}\in A,{\bm{Z}}_{k}\in B,Z_{k+1}\in C\mid{\bm{Y}}_{k})=\int_{A}\int_{B\times C}{f_{k+1}(y_{k+1},{\bm{z}}_{k+1}\mid{\bm{Y}}_{k})}\,\mathfrak{n}_{k+1}(d{\bm{z}}_{k+1})\,\widetilde{\mathfrak{n}}(dy_{k+1}).$
Hence,
$\int_{B\times C}{f_{k+1}(y_{k+1},{\bm{z}}_{k+1}\mid{\bm{Y}}_{k})}\,\mathfrak{n}_{k+1}(d{\bm{z}}_{k+1})\\ =\int_{B}\int_{C}{f_{k}(y_{k+1}\mid{\bm{Y}}_{k},{\bm{z}}_{k})\,\gamma_{k+1}({\bm{Y}}_{k},y_{k+1},{\bm{z}}_{k},dz_{k+1})f_{k}({\bm{z}}_{k}\mid{\bm{Y}}_{k})\,\mathfrak{n}_{k}(d{\bm{z}}_{k})},$
for $\widetilde{\mathfrak{n}}$-a.e. $y_{k+1}\in T$. In particular, with probability one, we have
$\mathcal{L}({\bm{Z}}_{k+1}\mid{\bm{Y}}_{k+1};d$ ${\bm{z}}_{k+1})$
$={f_{k+1}({\bm{z}}_{k+1}\mid{\bm{Y}}_{k+1})}\,\mathfrak{n}_{k+1}(d{\bm{z}}_{k+1})$
$=\frac{f_{k+1}(Y_{k+1},{\bm{z}}_{k+1}\mid{\bm{Y}}_{k})}{f_{k+1}(Y_{k+1}\mid{\bm{Y}}_{k})}\,\mathfrak{n}_{k+1}(d{\bm{z}}_{k+1})$
$=\frac{f_{k+1}({\bm{Y}}_{k})}{f_{k+1}({\bm{Y}}_{k+1})}{f_{k}(Y_{k+1}\mid{\bm{Y}}_{k},{\bm{z}}_{k})\,\gamma_{k+1}({\bm{Y}}_{k+1},{\bm{z}}_{k},dz_{k+1})f_{k}({\bm{z}}_{k}\mid{\bm{Y}}_{k})\,\mathfrak{n}_{k}(dz_{k})}.$
Since $f_{k+1}({\bm{y}}_{k})$ and $f_{k}({\bm{y}}_{k})$ are both densities of ${\bm{Y}}_{k}$ with respect to $\widetilde{\mathfrak{n}}^{k}$, we have $f_{k+1}({\bm{y}}_{k})=f_{k}({\bm{y}}_{k})$, $\widetilde{\mathfrak{n}}^{k}$-a.e. In particular, $f_{k+1}({\bm{Y}}_{k})=f_{k}({\bm{Y}}_{k})$ a.s. Thus,
$f_{k+1}({\bm{z}}_{k+1}\mid{\bm{Y}}_{k+1})\,\mathfrak{n}_{k+1}(d{\bm{z}}_{k+1})\\ =\frac{f_{k}({\bm{Y}}_{k})}{f_{k+1}({\bm{Y}}_{k+1})}{f_{k}(Y_{k+1}\mid{\bm{Y}}_{k},{\bm{z}}_{k})\,\gamma_{k+1}^{*}(Y,{\bm{z}}_{k},dz_{k+1})f_{k}({\bm{z}}_{k}\mid{\bm{Y}}_{k})\,\mathfrak{n}_{k}(dz_{k})},$
almost surely. Note that $\gamma_{1}^{*}(Y,dz_{1})=\gamma_{1}(Y_{1},dz_{1})=f_{1}(z_{1}\mid Y_{1})\,\mathfrak{n}_{1}(dz_{1})$. Hence, starting with $k=M-1$ and iterating backwards to $k=1$, we obtain
$\mathcal{L}(Z\mid Y;dz)={\frac{1}{f_{M}(Z)}\,w(Y,Z)(\gamma_{1}^{*}\cdots\gamma_{M}^{*})(Y,dz)}.$
Since $\mathfrak{m}^{*}=\gamma_{1}^{*}\cdots\gamma_{M}^{*}$, this proves (4.7). ∎
5 Sequential imputation for IDPs
In this section, we apply sequential imputation, in the form of Theorem 4.7, to an array of samples from an IDP. In Theorem 4.7, we take $Z=\bm{\mu}_{M}$ and we let $Y$ represent some observations we have made about the samples ${\bm{X}}_{MN}$.
In Section 5.1, we show that sequential imputation cannot, in general, be used on direct observations $Y={\bm{X}}_{MN}$. Starting in Section 5.2, therefore, we assume that $Y$ is a discrete function of ${\bm{X}}_{MN}$. Section 5.2 contains our main result, Theorem 5.2. This theorem shows how to use sequential imputation to compute $\mathcal{L}(\bm{\mu}_{M}\mid Y)$. The chief challenge is to construct the simulated row distributions, $\bm{\mu}_{M}^{*,k}$. According to (4.5), these should be constructed using the single-row conditional distributions, $\mathcal{L}({\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}})$. In Section 5.3, we compute $\mathcal{L}({\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}})$. In Section 5.4, we use these to generate $\bm{\mu}_{M}^{*,k}$. Then, in Section 5.5, we prove Theorem 5.2.
5.1 The impossibility of direct observations
A direct observation of the samples would be represented by taking $Y={\bm{X}}_{MN}$. In general, though, we cannot treat the case $Y={\bm{X}}_{MN}$ with sequential imputation. This is because Assumption 4.6 may fail. More specifically, if $Y={\bm{X}}_{MN}$ and the base measure $\varrho$ is not discrete, then Assumption 4.6 will fail, according to Proposition 5.1 below.
To prove Proposition 5.1, we begin by establishing two formulas that we will need later. Let $A\in\mathcal{S}^{n}$ and $B\in\mathcal{M}_{1}$. By Theorem 2.3,
$P(\mu_{m}\in B,X_{mn}\in A\mid\bm{\mu}_{m-1})$ $=E[1_{B}(\mu_{m})P(X_{mn}\in A\mid\bm{\mu}_{m})\mid\bm{\mu}_{m-1}]$
$=E[1_{B}(\mu_{m})\mu_{m}^{n}(A)\mid\bm{\mu}_{m-1}].$
By (3.6), this gives
$P(\mu_{m}\in B,X_{mn}\in A\mid\bm{\mu}_{m-1})=\frac{1}{\kappa+m-1}\bigg{(}\kappa E[1_{B}(\mu_{m})\mu_{m}^{n}(A)]+\sum_{i=1}^{m-1}1_{B}(\mu_{i})\mu_{i}^{n}(A)\bigg{)}.$ (5.1)
In particular, since $P(X_{mn}\in A)=E[\mu_{m}^{n}(A)]$, we have
$P(X_{mn}\in A\mid\bm{\mu}_{m-1})=\frac{1}{\kappa+m-1}\bigg{(}\kappa P(X_{mn}\in A)+\sum_{i=1}^{m-1}\mu_{i}^{n}(A)\bigg{)}.$ (5.2)
Proposition 5.1.
Suppose $\varrho$ is not discrete. That is, there exists $B\in S$ such that $\varrho(B)>0$ and $\varrho(\{x\})=0$ for all $x\in B$. Then there do not exist $\sigma$-finite measures $\widetilde{\mathfrak{n}}_{1}$ and $\widetilde{\mathfrak{n}}_{2}$ on $S$ such that $\mathcal{L}(\xi_{11},\xi_{21})\ll\widetilde{\mathfrak{n}}_{1}\times\widetilde{\mathfrak{n}}_{2}$.
Proof.
Assume $\mathcal{L}(\xi_{11},\xi_{21})\ll\widetilde{\mathfrak{n}}_{1}\times\widetilde{\mathfrak{n}}_{2}$. Let $A=\{x\in S:\widetilde{\mathfrak{n}}_{1}(\{x\})>0\}$, so that $A$ is countable. Let $D=\{(x,x)\in S^{2}:x\in B\cap A^{c}\}$. Then
$(\widetilde{\mathfrak{n}}_{1}\times\widetilde{\mathfrak{n}}_{2})(D)=\int_{B\cap A^{c}}\widetilde{\mathfrak{n}}_{1}(\{x\})\,\widetilde{\mathfrak{n}}_{2}(dx)=0.$
Therefore, by hypothesis,
$P(\xi_{21}=\xi_{11},\xi_{11}\in B\cap A^{c})=P((\xi_{11},\xi_{21})\in D)=0.$ (5.3)
On the other hand, by Theorem 2.5 and (5.2),
$P(\xi_{21}\in C\mid\xi_{11})$ $=E[P(\xi_{21}\in C\mid\mu_{1},\xi_{11})\mid\xi_{11}]$
$=E[P(\xi_{21}\in C\mid\mu_{1})\mid\xi_{11}]$
$=E\bigg{[}\frac{1}{\kappa+1}(\kappa P(\xi_{21}\in C)+\mu_{1}(C))\;\bigg{|}\;\xi_{11}\bigg{]}$
$=\frac{\kappa}{\kappa+1}\,\varrho(C)+\frac{1}{\kappa+1}E[\mu_{1}(C)\mid\xi_{11}].$
By (3.5) and Theorem 3.1, this gives
$P(\xi_{21}\in C\mid\xi_{11})=\frac{\kappa}{\kappa+1}\,\varrho(C)+\frac{1}{\kappa+1}\left(\frac{\varepsilon}{\varepsilon+1}\,\varrho(C)+\frac{1}{\varepsilon+1}\,\delta_{\xi_{11}}(C)\right).$
Hence,
$\mathcal{L}(\xi_{21}\mid\xi_{11})=\frac{\tau}{\tau+1}\,\varrho+\frac{1}{\tau+1}\,\delta_{\xi_{11}},$
where $\tau=(\kappa+1)(\varepsilon+1)-1$. We therefore have
$P(\xi_{21}=\xi_{11},\xi_{11}\in B\cap A^{c})$ $=E[1_{B\cap A^{c}}(\xi_{11})P(\xi_{21}=\xi_{11}\mid\xi_{11})]$
$=E\bigg{[}1_{B\cap A^{c}}(\xi_{11})\bigg{(}\frac{\tau}{\tau+1}\,\varrho(\{\xi_{11}\})+\frac{1}{\tau+1}\bigg{)}\bigg{]}.$
Note that $\xi_{11}\in B$ implies $\varrho(\{\xi_{11}\})=0$. Thus,
$P(\xi_{21}=\xi_{11},\xi_{11}\in B\cap A^{c})=\frac{1}{\tau+1}P(\xi_{11}\in B\cap A^{c})=\frac{1}{\tau+1}\,\varrho(B\cap A^{c}).$
Since $A$ is countable,
$\varrho(B\cap A^{c})=\varrho(B)-\sum_{x\in A\cap B}\varrho(\{x\})=\varrho(B)>0,$
which contradicts (5.3). ∎
5.2 Sequential imputation with discrete observations
If $\varrho$ is discrete, then we may assume without loss of generality that $S$ is countable. Therefore, if $S$ is uncountable and $Y_{m}=X_{mN}=(\xi_{m1},\ldots,\xi_{mN})$, then Proposition 5.1 shows that Assumption 4.6 fails. In other words, if $S$ is uncountable and we wish to use sequential imputation, then we cannot observe the data $\xi_{ij}$ directly.
On the other hand, we can observe discrete functions of $\xi_{ij}$. This is because Assumption 4.6 is trivially satisfied whenever $Y$ is discrete. From an applied perspective, this is no restriction at all. Any real-world measurement will have limits to its precision, meaning that only a finite number of measurement outcomes are possible.
Theorem 5.2 below describes how to use sequential imputation to compute $\mathcal{L}(\bm{\mu}_{M}\mid Y)$ when our observations $Y$ are discrete. The proof of Theorem 5.2 will be given in Section 5.5. Before stating Theorem 5.2, we first establish some notation.
Let $T$ be a countable set, fix $N\in\mathbb{N}$, and let $\varphi_{m}:S^{N}\to T$. Define $Y_{m}=\varphi_{m}(X_{mN})$. We adopt the notation of Section 4.3, so that $Y=(Y_{1},\ldots,Y_{M})$, ${\bm{Y}}_{m}=(Y_{1},\ldots,Y_{m})$, $y=(y_{1},\ldots,y_{M})\in T^{M}$, and ${\bm{y}}_{m}=(y_{1},\ldots,y_{m})\in T^{m}$. We will apply Theorem 4.7 with $M_{1}$ in place of $S$ and $\bm{\mu}_{M}$ in place of $Z$. We therefore change notation from $z$ to $\nu$. That is, $\nu=(\nu_{1},\ldots,\nu_{M})\in M_{1}^{M}$ and $\bm{\nu}_{m}=(\nu_{1},\ldots,\nu_{m})\in M_{1}^{m}$.
Let $\varrho_{n}=\mathcal{L}(X_{mn})$, so that $\varrho_{1}=\varrho$. Using (3.8) with $\varepsilon\varrho$ instead of $\kappa\rho$ gives us a recursive way to compute $\varrho_{n}$. In particular, for $B_{n}\in\mathcal{S}^{n}$ and $B\in\mathcal{S}$, we have
$\varrho_{n+1}(B_{n}\times B)=\frac{\varepsilon}{\varepsilon+n}\varrho_{n}(B_{n})\varrho(B)+\frac{1}{\varepsilon+n}\sum_{i=1}^{n}\varrho_{n}(B_{n}\cap\pi_{i}^{-1}B),$
where $\pi_{i}:S^{n}\to S$ is the projection onto the $i$th co-ordinate.
For $m\in\{1,\ldots,M\}$ and $y_{m}\in T$, let $A_{m}=\varphi_{m}^{-1}(\{y_{m}\})\in\mathcal{S}^{N}$. Then $Y_{m}=y_{m}$ if and only if $X_{mN}\in A_{m}$. Therefore, the prior likelihoods, $P(Y_{m}=y_{m})$, satisfy
$P(Y_{m}=y_{m})=\varrho_{N}(A_{m}).$ (5.4)
Although the notation does not explicitly indicate it, we must remember that the set $A_{m}$ depends on the vector $y_{m}$.
Now fix $y=(y_{1},\ldots,y_{M})\in T^{M}$. Using $y$, we will construct a weighted simulation of $\bm{\mu}_{M}$, which is a pair $(t,\bm{u}_{M}^{*})$, where $t=\{t_{mi}:1\leq m\leq M,1\leq i\leq m\}$ is a triangular array of $[0,\infty)$-valued random variables and $\bm{u}_{M}^{*}=(u_{1}^{*},\ldots,u_{M}^{*})$ is a vector of $M_{1}$-valued random variables, all of which are independent of $Y$. The rows of $t$, which we denote by $t_{m}=(t_{m1},\ldots,t_{mm})$, are called the row weights of the weighted simulation, and the random measures $u_{m}^{*}$ are called the simulated row distributions. We construct $t_{m}$ and $u_{m}^{*}$ by recursion on $m$ as follows. Let
$t_{mi}=\begin{cases}(u_{i}^{*})^{N}(A_{m})&\text{if $1\leq i<m$},\\ \kappa\varrho_{N}(A_{m})&\text{if $i=m$},\end{cases}$ (5.5)
and
$\mathcal{L}(u_{m}^{*}\mid\bm{u}_{m-1}^{*})\propto t_{mm}\int_{S^{N}}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{n=1}^{N}\delta_{x_{n}}\bigg{)}}\,\varrho_{N}(dx\mid A_{m})+\sum_{i=1}^{m-1}t_{mi}\delta_{u_{i}^{*}},$ (5.6)
where $\bm{u}_{m-1}^{*}=(u_{1}^{*},\ldots,u_{m-1}^{*})$ and $\varrho_{N}(A\mid A_{m})=P(X_{mN}\in A\mid X_{mN}\in A_{m})$. In other words,
$P(u_{m}^{*}=u_{i}^{*}\mid\bm{u}_{m-1}^{*})={\frac{t_{mi}}{t_{m1}+\cdots t_{mm}}},$
for $1\leq i<m$, and, with probability $t_{mm}/(t_{m1}+\cdots+t_{mm})$, the random measure $u_{m}^{*}$ is independent of $\bm{u}_{m-1}^{*}$ and has distribution
$\int_{S^{N}}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{n=1}^{N}\delta_{x_{n}}\bigg{)}}\,\varrho_{N}(dx\mid A_{m}).$ (5.7)
Finally, having constructed the weighted simulation $(t,\bm{u}_{M}^{*})$, we define the total weight of the weighted simulation to be
$V=\prod_{m=1}^{M}\frac{1}{\kappa+m-1}\sum_{i=1}^{m}t_{mi}.$ (5.8)
Theorem 5.2.
Let $\{(t^{k},\bm{u}_{M}^{*,k})\}_{k=1}^{K}$ be $K$ independent weighted simulations as above, with corresponding total weights $V_{k}$. Then
$\mathcal{L}(\bm{\mu}_{M}\mid Y=y)=\lim_{K\to\infty}\frac{\sum_{k=1}^{K}V_{k}\delta(\bm{u}_{M}^{*,k})}{\sum_{k=1}^{K}V_{k}},$ (5.9)
where $\delta(\bm{u}_{M}^{*,k})$ is the point mass measure on $M_{1}^{M}$ centered at $\bm{u}_{M}^{*,k}$. Consequently, if $\Phi$ is a measurable function on $M_{1}^{M}$ taking values in a metric space and $P(\bm{\mu}_{M}\in D^{c}\mid Y=y)=1$, where $D\subseteq M_{1}^{M}$ is the set of discontinuities of $\Phi$, then
$\mathcal{L}(\Phi(\bm{\mu}_{M})\mid Y=y)=\lim_{K\to\infty}\frac{\sum_{k=1}^{K}V_{k}\delta(\Phi(\bm{u}_{M}^{*,k}))}{\sum_{k=1}^{K}V_{k}}.$ (5.10)
The proof of Theorem 5.2 will be given in Section 5.5.
Corollary 5.3.
With the assumptions of Theorem 5.2, we have
$\mathcal{L}(\mu_{M+1}\mid Y=y)=\lim_{K\to\infty}\frac{1}{\kappa+M}\left({\kappa\mathcal{D}(\varepsilon\varrho)+\sum_{m=1}^{M}\frac{\sum_{k=1}^{K}V_{k}\delta(u_{m}^{*,k})}{\sum_{k=1}^{K}V_{k}}}\right).$ (5.11)
Consequently, if $\Phi$ is a measurable function on $M_{1}$ taking values in a metric space and $P(\mu_{M+1}\in D^{c}\mid Y=y)=1$, where $D\subseteq M_{1}$ is the set of discontinuities of $\Phi$, then
$\mathcal{L}(\Phi(\mu_{M+1})\mid Y=y)=\lim_{K\to\infty}\frac{1}{\kappa+M}\left({\kappa\mathcal{D}(\varepsilon\varrho)\circ\Phi^{-1}+\sum_{m=1}^{M}\frac{\sum_{k=1}^{K}V_{k}\delta(\Phi(u_{m}^{*,k}))}{\sum_{k=1}^{K}V_{k}}}\right).$ (5.12)
Proof.
Let $\Psi:M_{1}\to\mathbb{R}$ be continuous and bounded. By Theorem 2.5 and (3.6), we have
$E[\Psi(\mu_{M+1})\mid Y]$ $=E[E[\Psi(\mu_{M+1})\mid\bm{\mu}_{M},Y]\mid Y]$
$=E[E[\Psi(\mu_{M+1})\mid\bm{\mu}_{M}]\mid Y]$
$=E\bigg{[}\frac{\kappa}{\kappa+M}\int_{M_{1}}\Psi(\nu)\,\mathcal{D}(\varepsilon\varrho,d\nu)+\frac{1}{\kappa+M}\sum_{m=1}^{M}\Psi(\mu_{m})\;\bigg{|}\;Y\bigg{]}$
$=\frac{\kappa}{\kappa+M}\int_{M_{1}}\Psi(\nu)\,\mathcal{D}(\varepsilon\varrho,d\nu)+\frac{1}{\kappa+M}\sum_{m=1}^{M}E[\Psi(\mu_{m})\mid Y]$
By (5.10), this gives
$E[\Psi(\mu_{M+1})\mid Y]=\frac{\kappa}{\kappa+M}\int_{M_{1}}\Psi(\nu)\,\mathcal{D}(\varepsilon\varrho,d\nu)+\frac{1}{\kappa+M}\sum_{m=1}^{M}\lim_{K\to\infty}{\frac{\sum_{k=1}^{K}V_{k}\Psi(u_{m}^{*,k})}{\sum_{k=1}^{K}V_{k}}}.$
Since $\Psi$ was arbitrary, this proves (5.11), and (5.12) follows immediately. ∎
5.3 Conditioning on a single row
We will prove Theorem 5.2 by applying Theorem 4.7. To do this, we must, among other things, compute the conditional distribution $\gamma_{m}$ described in Section 4.3. This is done below, and the result is presented in (5.14).
Theorem 5.4.
Fix $m\in\{1,\ldots,M\}$. Let $\gamma_{m}$ be the probability kernel from $T^{m}\times M_{1}^{m-1}$ to $M_{1}$ with ${\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}}\sim\gamma_{m}({\bm{Y}}_{m},\bm{\mu}_{m-1})$. Fix ${\bm{y}}_{m}\in T^{m}$ and $\bm{\nu}_{m-1}\in M_{1}^{m-1}$. For $1\leq i\leq m$, let
$q_{i}=q_{i}^{m}(\bm{\nu}_{m-1})=\begin{cases}\nu_{i}^{N}(A_{m})&\text{if $1\leq i<m$},\\ \kappa\varrho_{N}(A_{m})&\text{if $i=m$},\end{cases}$ (5.13)
and let $p_{i}=q_{i}/(q_{1}+\cdots+q_{m})$. Then
$\gamma_{m}({\bm{y}}_{m},\bm{\nu}_{m-1})=p_{m}\int_{S^{N}}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{j=1}^{N}\delta_{x_{j}}\bigg{)}}\,\varrho_{N}(dx\mid A_{m})+\sum_{i=1}^{m-1}p_{i}\delta_{\nu_{i}}.$ (5.14)
Proof.
Let $\gamma$ be the probability kernel on the right-hand side of (5.14) and let $B\in\mathcal{M}_{1}$. We must show that $P(\mu_{m}\in B\mid{\bm{Y}}_{m},\bm{\mu}_{m-1})=\gamma({\bm{Y}}_{m},\bm{\mu}_{m-1},B)$. By Theorem 2.5, it suffices to show that $P(\mu_{m}\in B\mid Y_{m},\bm{\mu}_{m-1})=\gamma({\bm{Y}}_{m},\bm{\mu}_{m-1},B)$.
Define the kernel $\vartheta_{m}$ from $T\times M_{1}^{m-1}$ to $M_{1}$ by
$\vartheta_{m}(y_{m},\bm{\nu}_{m-1},d\nu_{m})=\kappa\nu_{m}^{N}(A_{m})\,\mathcal{L}(\mu_{m};d\nu_{m})+\sum_{i=1}^{m-1}\nu_{m}^{N}(A_{m})\,\delta_{\nu_{i}}(d\nu_{m}).$ (5.15)
Then (5.1) gives
$P(\mu_{m}\in B,Y_{m}=y_{m}\mid\bm{\mu}_{m-1})=\frac{\vartheta(y_{m},\bm{\mu}_{m-1},B)}{\kappa+m-1}.$ (5.16)
Now let $C\in\sigma(Y_{m},\bm{\mu}_{m-1})$. Without loss of generality, we may assume that $C$ is of the form $C=\{Y_{m}\in D\}\cap\{\bm{\mu}_{m-1}\in F\}$ for some $D\subseteq T$ and $F\in\mathcal{M}_{1}^{m-1}$. Then (5.16) gives
$E[1_{B}(\mu_{m})1_{C}]$ $=P(\mu_{m}\in B,Y_{m}\in D,\bm{\mu}_{m-1}\in F)$
$=E\bigg{[}1_{F}(\bm{\mu}_{m-1})\sum_{y_{m}\in D}{P(\mu_{m}\in B,Y_{m}=y_{m}\mid\bm{\mu}_{m-1})}\bigg{]}$
$=E\bigg{[}1_{F}(\bm{\mu}_{m-1})\sum_{y_{m}\in D}{\frac{\vartheta(y_{m},\bm{\mu}_{m-1},B)}{\vartheta(y_{m},\bm{\mu}_{m-1},M_{1})}P(Y_{m}=y_{m}\mid\bm{\mu}_{m-1})}\bigg{]},$
where in the last line we have used (5.16) with $B=M_{1}$. Hence,
$E[1_{B}(\mu_{m})1_{C}]$ $=\sum_{y_{m}\in D}E\bigg{[}1_{F}(\bm{\mu}_{m-1}){\frac{\vartheta(y_{m},\bm{\mu}_{m-1},B)}{\vartheta(y_{m},\bm{\mu}_{m-1},M_{1})}E[1_{\{y_{m}\}}(Y_{m})\mid\bm{\mu}_{m-1}]}\bigg{]}$
$=\sum_{y_{m}\in D}E\bigg{[}E\bigg{[}1_{F}(\bm{\mu}_{m-1}){\frac{\vartheta(y_{m},\bm{\mu}_{m-1},B)}{\vartheta(y_{m},\bm{\mu}_{m-1},M_{1})}1_{\{y_{m}\}}(Y_{m})\;\bigg{|}\;\bm{\mu}_{m-1}}\bigg{]}\bigg{]}$
$=\sum_{y_{m}\in D}E\bigg{[}1_{F}(\bm{\mu}_{m-1}){\frac{\vartheta(y_{m},\bm{\mu}_{m-1},B)}{\vartheta(y_{m},\bm{\mu}_{m-1},M_{1})}1_{\{y_{m}\}}(Y_{m})}\bigg{]}.$
We can rewrite this is
$E[1_{B}(\mu_{m})1_{C}]$ $=\sum_{y_{m}\in D}E\bigg{[}1_{F}(\bm{\mu}_{m-1}){\frac{\vartheta(Y_{m},\bm{\mu}_{m-1},B)}{\vartheta(Y_{m},\bm{\mu}_{m-1},M_{1})}1_{\{y_{m}\}}(Y_{m})}\bigg{]}$
$=E\bigg{[}1_{F}(\bm{\mu}_{m-1}){\frac{\vartheta(Y_{m},\bm{\mu}_{m-1},B)}{\vartheta(Y_{m},\bm{\mu}_{m-1},M_{1})}\sum_{y_{m}\in D}1_{\{y_{m}\}}(Y_{m})}\bigg{]}$
$=E\bigg{[}1_{F}(\bm{\mu}_{m-1}){\frac{\vartheta(Y_{m},\bm{\mu}_{m-1},B)}{\vartheta(Y_{m},\bm{\mu}_{m-1},M_{1})}1_{D}(Y_{m})}\bigg{]}$
$=E\bigg{[}\frac{\vartheta(Y_{m},\bm{\mu}_{m-1},B)}{\vartheta(Y_{m},\bm{\mu}_{m-1},M_{1})}1_{C}\bigg{]}.$
Hence,
$P(\mu_{m}\in B\mid Y_{m},\bm{\mu}_{m-1})=\frac{\vartheta(Y_{m},\bm{\mu}_{m-1},B)}{\vartheta(Y_{m},\bm{\mu}_{m-1},M_{1})}.$
It remains to show that $\gamma({\bm{y}}_{m},\bm{\mu}_{m-1})=\vartheta(y_{m},\bm{\mu}_{m-1})/\vartheta(y_{m},\bm{\mu}_{m-1},M_{1})$. Note that
$P(\mu_{m}\in B,Y_{m}=y_{m})$ $=E[1_{B}(\mu_{m})P(X_{mN}\in A_{m}\mid\mu_{m})]$
$=E[1_{B}(\mu_{m})\mu_{m}^{N}(A_{m})]$
$=\int_{B}\nu_{m}^{N}(A_{m})\,\mathcal{L}(\mu_{m};d\nu_{m}),$
which shows that
$\mathcal{L}(\mu_{m}\mid Y_{m}=y_{m};d\nu_{m})=\frac{1}{P(Y_{m}=y_{m})}\,\nu_{m}^{N}(A_{m})\,\mathcal{L}(\mu_{m};d\nu_{m}).$
Thus, (5.15) becomes
$\vartheta(y_{m},\bm{\nu}_{m-1})=\kappa P(Y_{m}=y_{m})\mathcal{L}(\mu_{m}\mid Y_{m}=y_{m})+\sum_{i=1}^{m-1}\nu_{i}^{N}(A_{m})\delta_{\nu_{i}}.$
By (3.10),
$\mathcal{L}(\mu_{m}\mid Y_{m}=y_{m})=\int_{S^{N}}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{n=1}^{N}\delta_{x_{n}}\bigg{)}}\,\mathcal{L}(X_{mN}\mid Y_{m}=y_{m};dx)$
Since $\{Y_{m}=y_{m}\}=\{X_{mN}\in A_{m}\}$ and $\varrho_{N}=\mathcal{L}(X_{mN})$, we can combine these last two equations to arrive at
$\vartheta(y_{m},\bm{\nu}_{m-1})=\kappa\varrho_{N}(A_{m})\int_{S^{N}}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{n=1}^{N}\delta_{x_{n}}\bigg{)}}\,\varrho_{N}(dx\mid A_{m})+\sum_{i=1}^{m-1}{\nu_{i}^{N}(A_{m})\delta_{\nu_{i}}}.$
It therefore follows from (5.14) that $\gamma({\bm{y}}_{m},\bm{\mu}_{m-1})=\vartheta(y_{m},\bm{\mu}_{m-1})/\vartheta(y_{m},\bm{\mu}_{m-1},M_{1})$. ∎
5.4 Generating the simulations
Having computed $\gamma_{m}$ in Theorem 5.4, we can now compute the simulation measure $\mathfrak{m}^{*}$, described in Section 4.3. In (5.6) and (5.8), the terms $u_{m}^{*}$ and $V$ depend on $y$. To emphasize this dependence, we write $u_{m}^{*,k}=u_{m}^{*,k}(y)$ and $V_{k}=V_{k}(y)$. Define $\mu_{m}^{*,k}=u_{m}^{*,k}(Y)$. Note that $u_{m}^{*,k}(y)$ is independent of $Y$, whereas $\mu_{m}^{*,k}$ is not. The random measure $\mu_{m}^{*,k}$ is playing the role of $Z_{m}^{*,k}$ in Section 4.3. To show that we have constructed $\mu_{m}^{*,k}$ correctly, we must show that $\{\bm{\mu}_{M}^{*,k}\}_{k=1}^{\infty}\mid Y\sim\mathfrak{m}^{*}(Y)^{\infty}$. This is done below in Proposition 5.5.
To prove Proposition 5.5, we use the following explicit construction of $u_{m}^{*,k}(y)$. Define $H\subseteq\mathbb{R}^{m}$ by $H=[0,\infty)^{\infty}\setminus\{(0,\ldots,0)\}$. Let
$U=\{U_{mk}(t):1\leq m\leq M,1\leq k\leq K,t\in H\}$
be an independent collection of random variables, where $U_{mk}(t)$ takes values in $\{1,\ldots,m\}$ and satisfies $P(U_{mk}(t)=i)=t_{i}/(t_{1}+\cdots+t_{m})$. Let
$\lambda=\{\lambda_{mk}:1\leq m\leq M,1\leq k\leq K\}$
be an independent collection of random measures on $S$, where $\lambda_{mk}$ is a Dirichlet mixture satisfying
$\lambda_{mk}\sim\int_{S^{N}}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{n=1}^{N}\delta_{x_{n}}\bigg{)}}\,\varrho_{N}(dx\mid A_{m}).$ (5.17)
Assume $U$, $\lambda$, and $Y$ are independent.
Define $t_{m}^{k}=(t_{m1}^{k},\ldots,t_{mm}^{k})\in\mathbb{R}^{m}$ and $\theta(m,k)\in\{1,\ldots,m\}$ recursively as follows. Let $t_{11}^{k}=\kappa\varrho_{N}(A_{1})$ and $\theta(1,k)=1$. For $m>1$, let
$t_{mi}^{k}=\begin{cases}\lambda_{\theta(i,k),k}^{N}(A_{m})&\text{if $1\leq i<m$},\\ \kappa\varrho_{N}(A_{m})&\text{if $i=m$},\end{cases}$ (5.18)
and
$\theta(m,k)=\begin{cases}\theta(U_{mk}(t^{mk}),k)&\text{if $1\leq U_{mk}(t^{mk})<m$},\\ m&\text{if $U_{mk}(t^{mk})=m$}.\end{cases}$ (5.19)
With this construction, we may write $u_{m}^{*,k}(y)=\lambda_{\theta(m,k),k}$.
In the proof of Proposition 5.5, we also use the notation $\mathcal{F}\vee\mathcal{G}=\sigma(\mathcal{F}\cup\mathcal{G})$, whenever $\mathcal{F}$ and $\mathcal{G}$ are $\sigma$-algebras on a common set.
Proposition 5.5.
Let $\gamma_{m}$ be as in Theorem 5.4 and $\mathfrak{m}^{*}=\gamma_{1}^{*}\cdots\gamma_{M}^{*}$, where $\gamma_{m}^{*}(y,\bm{\nu}_{m-1})=\gamma_{m}({\bm{y}}_{m},\bm{\nu}_{m-1})$. Then $\{\bm{\mu}_{M}^{*,k}\}_{k=1}^{\infty}\mid Y\sim\mathfrak{m}^{*}(Y)^{\infty}$.
Proof.
As noted in (4.5), it suffices to show that $\mu_{m}^{*,k}\mid Y,\bm{\mu}_{m-1}^{*,k}\sim\gamma_{m}({\bm{Y}}_{m},\bm{\mu}_{m-1}^{*,k})$.
We first note that if $\mathcal{F}_{mk}=\sigma(U_{1k},\ldots,U_{mk},\lambda_{1k},\ldots,\lambda_{mk})$, then $t_{m}^{k}\in\mathcal{F}_{m-1,k}$ and $\theta(m,k)\in\mathcal{F}_{m-1,k}\vee\sigma(U_{mk})$. This follows from (5.18) and (5.19) by induction. Also, by (5.19), we have
$u_{m}^{*,k}(y)=\begin{cases}u_{i}^{*,k}(y)&\text{if $U_{mk}(t^{mk})=i<m$},\\ \lambda_{mk}&\text{if $U_{mk}(t^{mk})=m$}.\end{cases}$ (5.20)
Hence, $u_{m}^{*,k}(y)\in\mathcal{F}_{mk}$. In particular, $U_{mk}$, $\lambda_{mk}$, and $\bm{u}_{m-1}^{*,k}(y)$ are independent.
Now let $B\in\mathcal{M}_{1}$ and $C\in\sigma(Y,\bm{\mu}_{m-1}^{*,k})$. Without loss of generality, we may assume that $C$ is of the form $C=\{Y\in D\}\cap\{\bm{\mu}_{m-1}^{*,k}\in F\}$ for some $D\subseteq T^{M}$ and $F\in\mathcal{M}_{1}^{m-1}$. We then have
$E[1_{B}(\mu_{m}^{*,k})1_{C}]$ $=P(Y\in D,\bm{\mu}_{m}^{*,k}\in F\times B)$
$=\sum_{y\in D}P(Y=y,\bm{u}_{m}^{*,k}(y)\in F\times B)$
$=\sum_{y\in D}P(Y=y)P(\bm{u}_{m}^{*,k}(y)\in F\times B)$
$=\sum_{y\in D}P(Y=y)E[1_{F}(\bm{u}_{m-1}^{*,k}(y))P(u_{m}^{*,k}(y)\in B\mid\bm{u}_{m-1}^{*,k}(y))]$ (5.21)
Using (5.20), we obtain
$P(u_{m}^{*,k}(y)\in B\mid\bm{u}_{m-1}^{*,k}(y))=P(U_{mk}(t_{m}^{k})=m,\lambda_{mk}\in B\mid\bm{u}_{m-1}^{*,k}(y))\\ +\sum_{i=1}^{m-1}P(U_{mk}(t_{m}^{k})=i,u_{i}^{*,k}(y)\in B\mid\bm{u}_{m-1}^{*,k}(y)).$
From (5.18) and (5.13), it follows that $t_{mi}^{k}=q_{i}^{m}(\bm{u}_{m-1}^{*,k}(y))$. Since $U_{mk}$, $\lambda_{mk}$, and $\bm{u}_{m-1}^{*,k}(y)$ are independent, the above becomes
$P(u_{m}^{*,k}(y)\in B\mid\bm{u}_{m-1}^{*,k}(y))=p_{m}^{m}(\bm{u}_{m-1}^{*,k}(y))P(\lambda_{mk}\in B)+\sum_{i=1}^{m-1}{p_{i}^{m}(\bm{u}_{m-1}^{*,k}(y))\delta_{u_{i}^{*,k}(y)}(B)}.$
It follows from (5.17) and (5.14) that
$P(u_{m}^{*,k}(y)\in B\mid\bm{u}_{m-1}^{*,k}(y))=\gamma_{m}({\bm{y}}_{m},\bm{u}_{m-1}^{*,k}(y),B).$
Substituting this into (5.21) and noting that $\bm{u}_{M}^{*,k}(y)$ and $Y$ are independent, we have
$E[1_{B}(\mu_{m}^{*,k})1_{C}]$ $=\sum_{y\in D}P(Y=y){E[1_{F}(\bm{u}_{m-1}^{*,k}(y))\gamma_{m}({\bm{y}}_{m},\bm{u}_{m-1}^{*,k}(y),B)]}$
$=\sum_{y\in D}E[1_{\{y\}}(Y)1_{F}(\bm{u}_{m-1}^{*,k}(y))\gamma_{m}({\bm{y}}_{m},\bm{u}_{m-1}^{*,k}(y),B)]$
$=\sum_{y\in D}E[1_{\{y\}}(Y)1_{F}(\bm{\mu}_{m-1}^{*,k})\gamma_{m}({\bm{Y}}_{m},\bm{\mu}_{m-1}^{*,k},B)]$
$=E[1_{D}(Y)1_{F}(\bm{\mu}_{m-1}^{*,k})\gamma_{m}({\bm{Y}}_{m},\bm{\mu}_{m-1}^{*,k},B)]$
$=E[\gamma_{m}({\bm{Y}}_{m},\bm{\mu}_{m-1}^{*,k},B)1_{C}],$
showing that $P(\mu_{m}^{*,k}\in B\mid Y,\bm{\mu}_{m-1}^{*,k})=\gamma_{m}({\bm{Y}}_{m},\bm{\mu}_{m-1}^{*,k},B)$. ∎
5.5 Proof of the main result
Having established Theorem 5.4 and Proposition 5.5, we are now ready to prove the main result.
Proof of Theorem 5.2.
We apply Theorem 4.7. Let $\gamma_{m}$ and $\mathfrak{m}^{*}$ be as in Proposition 5.5.
If $\widetilde{\mathfrak{n}}$ is counting measure on $T$ and $\mathfrak{n}_{m}=\mathcal{L}(\bm{\mu}_{m})$, then $\mathcal{L}(Y,\bm{\mu}_{m})\ll\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}_{m}$, so that Assumption 4.6 holds. Let $f_{m}$ be the density of $(Y,\bm{\mu}_{m})$ with respect to $\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}_{m}$, and recall the notational conventions of Section 4.4. Let $w(y,\nu)$ be given by (4.6).
Let $\bm{\mu}_{M}^{*,k}$ be as in Proposition 5.5, so that $\{\bm{\mu}_{M}^{*,k}\}_{k=1}^{\infty}\mid Y\sim\mathfrak{m}^{*}(Y)^{\infty}$. We define the weights $W_{k}=w(Y,\bm{\mu}_{M}^{*,k})$. We first prove that $W_{k}=V_{k}(Y)$, where, according to (5.8), we have
$V_{k}(y)=\prod_{m=1}^{M}\frac{1}{\kappa+m-1}\sum_{i=1}^{m}t_{mi}^{k}.$ (5.22)
Note that
$f_{1}(y_{1})=P(Y_{1}=y_{1})=\varrho_{N}(A_{1}).$
For the other factors in (4.6), we use Theorem 2.5, (5.2), and (5.4) to obtain
$P(Y_{m+1}=y_{m+1}\mid{\bm{Y}}_{m},\bm{\mu}_{m})$ $=P(X_{m+1,N}\in A_{m+1}\mid\bm{\mu}_{m})$
$=\frac{1}{\kappa+m}\bigg{(}\kappa\varrho_{N}(A_{m+1})+\sum_{i=1}^{m}\mu_{i}^{N}(A_{m+1})\bigg{)},$
so that
$f_{m}(y_{m+1}\mid{\bm{y}}_{m},\bm{\nu}_{m})=\frac{\kappa}{\kappa+m}\varrho_{N}(A_{m+1})+\frac{1}{\kappa+m}\sum_{i=1}^{m}\nu_{i}^{N}(A_{m+1}).$
Substituting this into (4.6) gives
$w(y,\nu)=\varrho_{N}(A_{1})\prod_{m=1}^{M-1}\bigg{(}\frac{\kappa}{\kappa+m}\varrho_{N}(A_{m+1})+\frac{1}{\kappa+m}\sum_{i=1}^{m}\nu_{i}^{N}(A_{m+1})\bigg{)},$
which can be rewritten as
$w(y,\nu)=\prod_{m=1}^{M}\bigg{(}\frac{\kappa}{\kappa+m-1}\varrho_{N}(A_{m})+\frac{1}{\kappa+m-1}\sum_{i=1}^{m-1}\nu_{i}^{N}(A_{m})\bigg{)}.$ (5.23)
In the proof of Proposition 5.5, we noted that $t_{mi}^{k}=q_{i}^{m}(\bm{u}_{m-1}^{*,k}(y))$. Hence, by (5.22) and (5.13), we have
$V_{k}(y)$ $=\prod_{m=1}^{M}{\frac{1}{\kappa+m-1}\sum_{i=1}^{m}q_{i}^{m}(\bm{u}_{m-1}^{*,k}(y))}$
$=\prod_{m=1}^{M}{\frac{1}{\kappa+m-1}\bigg{(}\kappa\varrho_{N}(A_{m})+\sum_{i=1}^{m-1}(u_{i}^{*,k}(y))^{N}(A_{m})\bigg{)}}.$
It follows from (5.23) that $V_{k}(y)=w(y,\bm{u}_{M}^{*,k}(y))$, so that $W_{k}=V_{k}(Y)$.
Finally, we construct $\widetilde{\bm{\mu}}_{M}^{K}$ so that
$\widetilde{\bm{\mu}}_{M}^{K}\mid\bm{\mu}_{M}^{*,1},\ldots,\bm{\mu}_{M}^{*,K},Y\propto\sum_{k=1}^{K}W_{k}\delta(\bm{\mu}_{M}^{*,k}).$ (5.24)
Since $f_{M}(y)=P(Y=y)$, we have
$\int_{T^{M}}f_{M}(y)^{2}\,\widetilde{\mathfrak{n}}^{M}(dy)=\sum_{y\in T^{M}}P(Y=y)^{2}\leq\sum_{y\in T^{M}}P(Y=y)=1,$
so that $f_{M}(y)\in L^{2}(\widetilde{\mathfrak{n}}^{M})$. Hence, by Theorem 4.7,
$\mathcal{L}(\bm{\mu}_{M}\mid Y)=\lim_{K\to\infty}{\mathcal{L}(\widetilde{\bm{\mu}}_{M}^{K}\mid\bm{\mu}_{M}^{*,1},\ldots,\bm{\mu}_{M}^{*,K},Y)}.$
Applying (5.24) to the above gives
$\mathcal{L}(\bm{\mu}_{M}\mid Y)=\lim_{K\to\infty}\frac{\sum_{k=1}^{K}W_{k}\delta(\bm{\mu}_{M}^{*,k})}{\sum_{k=1}^{K}W_{k}}.$
Since $W_{k}=V_{k}(Y)$ and $\bm{\mu}_{M}^{*,k}=\bm{u}_{M}^{*,k}(Y)$, this proves (5.9), and (5.10) follows immediately. ∎
6 Examples
In this section, we present four hypothetical applications to illustrate the use of the IDP. See https://github.com/jason-swanson/idp for the code used to generate the simulations.
The framework for each of these examples was described in Section 1. Namely, we consider a situation in which there are several agents, all from the same population. Each agent undertakes a sequence of actions. These actions are chosen according to the agent’s particular tendencies. Although different agents have different tendencies, there may be patterns in the population.
We observe a certain set of agents over a certain amount of time. Based on these observations, we want to make probabilistic forecasts about two things:
  • The future behavior of the agents we have observed.
  • The behavior of a new (unobserved) agent from the population.
We model this situation with an IDP. More specifically, let $S$ be a complete and separable metric space. Let $\xi$ be a row exchangeable array of $S$-valued random variables whose row distribution generator—the random measure $\varpi$ in Theorem 2.2—is an IDP. That is, $\varpi\sim\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$, where $\kappa>0$ and $\varepsilon>0$ are the column and row concentration parameters, respectively, and $\varrho\in M_{1}$ is the base measure. The value of $\xi_{ij}$ represents the $j$th action of the $i$th agent. In this way, the space $S$ represents the set of possible actions that the agents may undertake.
All the examples in this section involve a finite state space $S$. But, as we describe in Remark 6.2, this special case is easily generalized to the case of an arbitrary $S$ in which our observations are made with limited precision.
The outline of this section is as follows. In Section 6.1, we describe how Theorem 5.2 and Corollary 5.3 simplify in the case that $S$ is finite. After that, the remainder of the section is devoted to the examples. Our simplest example is in Section 6.2, and concerns the malfunctioning pressed penny machine described in Section 1. Section 6.3 presents a similar example, but with significantly more data. This is the same example treated in [12] (originally considered in [2]) and is concerned with the flicking of thumbtacks. Section 6.4 applies the IDP model to the analysis of Amazon reviews. The final example, found in Section 6.6, is about video game leaderboards. To prepare for that example, a custom prior distribution, which we call the “gamer” distribution, is presented in Section 6.5.
6.1 IDPs on a finite state space
Let $L\geq 2$ be an integer and suppose that $S=\{0,\ldots,L-1\}$. Let $p_{\ell}=\varrho(\{\ell\})$, so that we may identify $\varrho$ with the vector $\bm{p}=(p_{0},\ldots,p_{L-1})$. We assume that $p_{\ell}>0$ for all $\ell\in S$.
Let $y=\{y_{mn}:1\leq m\leq M,1\leq n\leq N_{m}\}$ be a jagged array of elements in $S$. The array $y$ denotes our observed data. That is, we observe $\xi_{mn}=y_{mn}$ for $1\leq m\leq M$ and $1\leq n\leq N_{m}$, and we wish to compute the conditional distribution of $\xi$ given these observations. Define the row counts $\overline{y}=\{\overline{y}_{m\ell}:1\leq m\leq M,0\leq\ell\leq L-1\}$ by $\overline{y}_{m\ell}=|\{n:y_{mn}=\ell\}|$. Since $\xi$ is row exchangeable, all of our calculations will depend on $y$ only through the array $\overline{y}$. We use $\overline{y}_{m}$ to denote the vector $(\overline{y}_{m1},\ldots,\overline{y}_{m,L-1})$.
To apply Theorem 5.2, let $N=\max\{N_{1},\ldots,N_{M}\}$ and $T=\bigcup_{n=1}^{N}S^{n}$. Let $\varphi_{m}:S^{N}\to T$ be the projection onto the first $N_{m}$ components, so that $\varphi_{m}(x_{1},\ldots,x_{N})=(x_{1},\ldots,x_{N_{m}})$. Then $Y=(Y_{1},\ldots,Y_{M})$, where $Y_{m}=\varphi_{m}(X_{mN})=X_{mN_{m}}$. Note that $A_{m}=\{Y_{m}=y_{m}\}=\{X_{mN_{m}}=y_{m}\}$. Therefore, if we define $\theta_{m\ell}=\mu_{m}(\{\ell\})$, then the prior likelihoods satisfy
$\varrho_{N}(A_{m})=P(X_{mN_{m}}=y_{m})=E[P(X_{mN_{m}}=y_{m}\mid\mu_{m})]=E\bigg{[}\prod_{\ell=0}^{L-1}\theta_{m\ell}^{\overline{y}_{m\ell}}\bigg{]}.$
From (3.1) it follows that
${(\theta_{m0},\ldots,\theta_{m,L-1})\sim\operatorname{Dir}(\varepsilon p_{0},\ldots,\varepsilon p_{L-1}).}$
This gives
$\varrho_{N}(A_{m})=E\bigg{[}\prod_{\ell=0}^{L-1}\theta_{m\ell}^{\overline{y}_{m\ell}}\bigg{]}=\frac{1}{B(\varepsilon\bm{p})}\int_{\Delta^{L-1}}{\prod_{\ell=0}^{L-1}t_{\ell}^{\overline{y}_{m\ell}+\varepsilon p_{\ell}-1}}\,dt=\frac{B(\varepsilon\bm{p}+\overline{y}_{m})}{B(\varepsilon\bm{p})},$ (6.1)
where $B({\bm{x}})=\Gamma(\sum_{\ell=0}^{L}x_{\ell})^{-1}\prod_{\ell=0}^{L}\Gamma(x_{\ell})$ is the multivariate Beta function.
Having computed the prior likelihoods, we turn our attention to the weighted simulations. From (3.10), it follows that (5.7) is equal to $\mathcal{L}(\mu_{m}\mid Y_{m}=y_{m})$. But $Y_{m}=X_{mN_{m}}$, so by (3.5) we can rewrite (5.5) and (5.6) as
$t_{mi}=\begin{cases}\prod_{n=1}^{N_{m}}u_{i}^{*}(y_{mn})&\text{if $1\leq i<m$},\\ \kappa\varrho_{N}(A_{m})&\text{if $i=m$},\end{cases}$
and
$\mathcal{L}(u_{m}^{*}\mid\bm{u}_{m-1}^{*})\propto t_{mm}{\mathcal{D}\bigg{(}\varepsilon\varrho+\sum_{n=1}^{N_{m}}\delta_{y_{mn}}\bigg{)}}+\sum_{i=1}^{m-1}t_{mi}\delta_{u_{i}^{*}}.$ (6.2)
If we define $\theta_{m\ell}^{*}=u_{m}^{*}(\{\ell\})$, then we can rewrite the row weights as
$t_{mi}=\begin{cases}\prod_{\ell=0}^{L-1}(\theta_{i\ell}^{*})^{\overline{y}_{m\ell}}&\text{if $1\leq i<m$},\\ \kappa\varrho_{N}(A_{m})&\text{if $i=m$}.\end{cases}$
In this case, we can identify $u_{m}^{*}$ with the vector $\theta_{m}^{*}=(\theta_{m0}^{*},\ldots,\theta_{m,L-1}^{*})$, and (6.2) becomes
$\mathcal{L}(\theta_{m}^{*}\mid\bm{\theta}_{m-1}^{*})\propto t_{mm}\operatorname{Dir}(\varepsilon\bm{p}+\overline{y}_{m})+\sum_{i=1}^{m-1}t_{mi}\delta_{\theta_{i}^{*}},$
where $\bm{\theta}_{m}^{*}=(\theta_{1}^{*},\ldots,\theta_{m}^{*})$. In other words, for $i<m$, we have $\theta_{m}^{*}=\theta_{i}^{*}$ with probabilty $t_{mi}$, and, with probability $t_{mm}/(t_{m1}+\cdots+t_{mm})$, the random vector $\theta_{m}^{*}$ is independent of $\bm{\theta}_{m-1}^{*}$ and has the Dirichlet distribution, $\operatorname{Dir}(\varepsilon\bm{p}+\overline{y}_{m})$.
Finally, we define $V$, the total weight of the simulation. According to (5.8), the total weight should be
$\prod_{m=1}^{M}\frac{1}{\kappa+m-1}\sum_{i=1}^{m}t_{mi}.$ (6.3)
But in Theorem 5.2, we see that the weights are all relative to their sum, so we are free to multiply this value by any constant that does not depend on $k$. Leaving it as it is will produce a very small number, on the order of $1/M!$. For computational purposes, then, we multiply (6.3) by $c^{M}M!$, where $c$ is a nonrandom constant. The total weight of our simulation is then
$V=\prod_{m=1}^{M}\frac{cm}{\kappa+m-1}\sum_{i=1}^{m}t_{mi}$ (6.4)
We call $\log c$ the log scale factor of the simulation. In the examples covered later in this section, we used $c=1$ unless otherwise specified.
Now, if $\theta=(\theta_{m\ell})\in\mathbb{R}^{M\times L}$ and $\Phi:\mathbb{R}^{M\times L}\to\mathbb{R}$ is continuous, then (5.10) gives
$\mathcal{L}(\Phi(\theta)\mid Y=y)=\lim_{K\to\infty}\frac{\sum_{k=1}^{K}V_{k}\delta(\Phi(\bm{\theta}_{M}^{*,k}))}{\sum_{k=1}^{K}V_{k}}.$ (6.5)
Similarly, if $\Phi:\mathbb{R}^{L}\to\mathbb{R}$ is continuous, then (5.12) gives
$\mathcal{L}(\Phi(\theta_{M+1})\mid Y=y)=\lim_{K\to\infty}\frac{1}{\kappa+M}\left({\kappa\operatorname{Dir}(\varepsilon\bm{p})\circ\Phi^{-1}+\sum_{m=1}^{M}\frac{\sum_{k=1}^{K}V_{k}\delta(\Phi(\theta_{m}^{*,k}))}{\sum_{k=1}^{K}V_{k}}}\right).$ (6.6)
Remark 6.1.
In the case $S=\{0,1\}$, the base measure $\varrho$ is entirely determined by the number $p=\varrho(\{1\})$, and we may define a single row count for each row, $\overline{y}_{m}=|\{n:y_{mn}=1\}|$. In this case, letting $a=\varepsilon p$ and $b=\varepsilon(1-p)$, we can rewrite (6.1) as
$\varrho_{N}(A_{m})=\frac{B(a+\overline{y}_{m},b+N_{m}-\overline{y}_{m})}{B(a,b)}.$
Defining $\theta_{m}^{*}=u_{m}^{*}(\{1\})$, the row weights of the weighted simulations become
$t_{mi}=\begin{cases}(\theta_{i}^{*})^{\overline{y}_{m}}(1-\theta_{i}^{*})^{N_{m}-\overline{y}_{m}}&\text{if $1\leq i<m$},\\ \kappa\varrho_{N}(A_{m})&\text{if $i=m$},\end{cases}$
and (6.2) becomes
$\mathcal{L}(\theta_{m}^{*}\mid\bm{\theta}_{m-1}^{*})\propto t_{mm}\operatorname{Beta}(a+\overline{y}_{m},b+N_{m}-\overline{y}_{m})+\sum_{i=1}^{m-1}t_{mi}\delta_{\theta_{i}^{*}}.$
Remark 6.2.
Let us return for the moment to the general setting, where $S$ is an arbitrary complete and separable metric space. Let $S^{\prime}$ be another complete and separable metric space and let $\psi:S\to S^{\prime}$ be measurable. Let $\xi^{\prime}=\{\xi^{\prime}_{ij}\}$, where $\xi^{\prime}_{ij}=\psi(\xi_{ij})$. It is straightforward to verify that $\xi^{\prime}$ is a row exchangeable array of $S^{\prime}$-valued random variables whose row distribution generator $\varpi^{\prime}$ satisfies $\varpi^{\prime}\sim\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho^{\prime}))$, where $\varrho^{\prime}=\varrho\circ\psi^{-1}$.
We can apply this to $S^{\prime}=\{0,\ldots,L-1\}$, where $L\geq 2$. For each $\ell\in S^{\prime}=\{0,\ldots,L-1\}$, choose $B_{\ell}\in\mathcal{S}$ so that $\{B_{\ell}:\ell\in S^{\prime}\}$ is a partition of $S$. Define $\psi:S\to S^{\prime}$ by $\psi=\sum_{\ell=0}^{L-1}\ell 1_{B_{\ell}}$ and let $\xi^{\prime}=\{\xi^{\prime}_{ij}\}$ where $\xi^{\prime}_{ij}=\psi(\xi_{ij})$. Suppose we can only observe the process $\xi^{\prime}$. That is, we can only observe the values of $\xi$ with enough precision to tell which piece of the partition those values lie in. Based on some set of these observations, we wish to make probabilistic inferences about $\xi^{\prime}$. Since $\xi^{\prime}$ is an array of samples from an IDP on $S^{\prime}$, we may do this using the simplified formulas in this section.
6.2 The pressed penny machine
Imagine a pressed penny machine, like those found in museums or tourist attractions. For a fee, the machine presses a penny into a commemorative souvenir. Now imagine the machine is broken, so that it mangles all the pennies we feed it. Each pressed penny it creates is mangled in its own way. Each has its own probability of landing on heads when flipped. In this situation, the agents are the pennies and the actions are the heads and tails that they produce.
Now suppose we create seven mangled pennies and flip each one 5 times, giving us the results in Table 1. Of the 35 flips, 23 of them (or about 65.7%) were heads. In fact, 6 of the 7 coins landed mostly on heads. The machine clearly seems predisposed to creating pennies that are biased towards heads.
Coin # 1st Flip 2nd Flip 3rd Flip 4th Flip 5th Flip
1 H H H H T
2 H T H H H
3 T H H T H
4 H H T H H
5 T T T H T
6 T H H H H
7 H T T H H
Table 1: Results of flipping seven different mangled pennies
Coin 5, though, produced only one head. Is this coin different from the others and actually biased toward tails? Or was it mere chance that its flips turned out that way? For instance, suppose all 7 coins had a 60% chance of landing on heads. In that case, there would still be a 43% chance that at least one of them would produce four tails. How should we balance these competing explanations and arrive at some concrete probabilities?
One way to answer this is to model the example with an IDP as in Section 6.1. We take $L=2$, so that $S=\{0,1\}$, where 00 represents tails and $1$ represents heads. We then take $\kappa=\varepsilon=1$ and $p_{0}=p_{1}=1/2$. From the table above, we have $M=7$, $N_{m}=5$ for all $m$, and
$y=\begin{pmatrix}1&1&1&1&0\\ 1&0&1&1&1\\ 0&1&1&0&1\\ 1&1&0&1&1\\ 0&0&0&1&0\\ 0&1&1&1&1\\ 1&0&0&1&1\end{pmatrix}.$
With $K=10000$, we generated the weighted simulations $(t^{k},\bm{\theta}_{7}^{*,k})$ for $1\leq k\leq K$, and computed their corresponding total weights, $V_{k}$. In this case, the effective sample size of our simulations (denoted by $K_{\varepsilon}^{\prime\prime}$ in Section 4.2) was approximately $6067$.
Before addressing Coin 5 directly, let us ask a different question. If we were to get a new coin from this machine, how would we expect it to behave? The new coin would have some random probability of heads, which is denoted by $\theta_{8,1}$. Taking $\Phi:\mathbb{R}^{2}\to\mathbb{R}$ to be the projection, $\Phi(x_{0},x_{1})=x_{1}$, we can use (6.6) to approximation the distribution of $\theta_{8,1}$, giving $\mathcal{L}(\theta_{8,1}\mid Y=y)\approx\nu$, where
$\nu=\frac{1}{8}\bigg{(}\operatorname{Beta}(1/2,1/2)+\sum_{m=1}^{7}\frac{\sum_{k=1}^{10000}V_{k}\delta(\theta_{m,1}^{*,k})}{\sum_{k=1}^{10000}V_{k}}\bigg{)}.$
Using this, we have $P(\xi_{8,1}=1\mid Y=y)=E[\theta_{8,1}\mid Y=y]\approx 0.633$, so that, given our observations, the first flip of a new coin has about a 63.3% chance of landing heads. To visualize the distribution of $\theta_{81}$ rather than simply its mean, we can plot the distribution function of $\nu$. See Figure 1(a) for a graph of $x\mapsto\nu((0,x])$.
(a) distribution function of $\theta_{8,1}$
(b) density of $\theta_{8,1}$ with $h\approx 0.119$
(c) density of $\theta_{8,1}$ with $h=0.001$
(d) density of $\theta_{5,1}$
Figure 1: Approximate distribution and density functions for $\theta_{m,\ell}$
For a different visualization, we can plot an approximate density for $\nu$. The measure $\nu$ has a discrete component, so we obtain an approximate density using Gaussian kernel density estimation, replacing each point mass $\delta_{x}$ by a Gaussian measure with mean $x$ and standard deviation $h$, where $h$ is the “bandwidth” of the density estimation. For the measure $\nu$, we used Python’s scipy.stats.gaussian_kde class to compute the bandwidth according to Scott’s Rule (see [13]). In this case, we obtained $h\approx 0.119$, yielding the graph in Figure 1(b). For a coarser estimate, see Figure 1(c), which uses $h=0.001$. In all the remaining examples in this section, we will default to using the bandwidth determined by Scott’s Rule.
Turning back to the question of Coin 5, if we define $\Phi:\mathbb{R}^{7\times 2}\to\mathbb{R}$ by $\Phi((x_{m,\ell}))=x_{5,1}$, then (6.5) gives
$\mathcal{L}(\theta_{5,1}\mid Y=y)\approx\frac{\sum_{k=1}^{10000}V_{k}\delta(\theta_{5,1}^{*,k})}{\sum_{k=1}^{10000}V_{k}}$
An approximate density for this measure is given in Figure 1(d). Using this, we can compute the probability that a sixth flip of Coin 5 lands on heads, which is
$P(\xi_{5,6}=1\mid Y=y)=E[\theta_{5,1}\mid Y=y]\approx 0.461.$
We can also compute the probability that Coin 5 is biased toward tails, which is given by $P(\theta_{51}<1/2\mid Y=y)\approx 0.481$.
6.3 Flicking thumbtacks
In [12], the following situation is considered. Imagine a box of 320 thumbtacks. We flick each thumbtack 9 times. If it lands point up, we call it a success. Point down is a failure. Because of the imperfections, each thumbtack has its own probability of success. The results (that is, the number of successes) for these 320 thumbtacks are given by
$r=($ $7,4,6,6,6,6,8,6,5,8,6,3,3,7,8,4,5,5,7,8,5,7,6,5,3,2,7,7,9,6,4,6,$
$4,7,3,7,6,6,6,5,6,6,5,6,5,6,7,9,9,5,6,4,6,4,7,6,8,7,7,2,7,7,4,6,$
$2,4,7,7,2,3,4,4,4,6,8,8,5,6,6,6,5,3,8,6,5,8,6,6,3,5,8,5,5,5,5,6,$
$3,6,8,6,6,6,8,5,6,4,6,8,7,8,9,4,4,4,4,6,7,1,5,6,7,2,3,4,7,5,6,5,$
$2,7,8,6,5,8,4,8,3,8,6,4,7,7,4,5,2,3,7,7,4,5,2,3,7,4,6,8,6,4,6,2,$
$4,4,7,7,6,6,6,8,7,4,4,8,9,4,4,3,6,7,7,5,5,8,5,5,5,6,9,1,7,3,3,5,$
$7,7,6,8,8,8,8,7,5,8,7,8,5,5,8,8,7,4,6,5,9,8,6,8,9,9,8,8,9,5,8,6,$
$3,5,9,8,8,7,6,8,5,9,7,6,5,8,5,8,4,8,8,7,7,5,4,2,4,5,9,8,8,5,7,7,$
$2,6,2,7,6,5,4,4,6,9,3,9,4,4,1,7,4,4,5,9,4,7,7,8,4,6,7,8,7,4,3,5,$
7,7,4,4,6,4,4,2,9,9,8,6,8,8,4,5,7,5,4,6,8,7,6,6,8,6,9,6,7,6,6,6) 7,7,4,4,6,4,4,2,9,9,8,6,8,8,4,5,7,5,4,6,8,7,6,6,8,6,9,6,7,6,6,6).7 , 7 , 4 , 4 , 6 , 4 , 4 , 2 , 9 , 9 , 8 , 6 , 8 , 8 , 4 , 5 , 7 , 5 , 4 , 6 , 8 , 7 , 6 , 6 , 8 , 6 , 9 , 6 , 7 , 6 , 6 , 6 ) .
This data originally came from an experiment described in [2]. In the original experiment, there were not 320 thumbtacks. Rather, there were 16 thumbtacks, 2 flickers, and 10 surfaces. We follow [12], however, in treating the data as if it came from 320 distinct thumbtacks.
To model this example we take $L=2$, so that $S=\{0,1\}$, where 00 represents failure (point down) and $1$ represents success (point up). To match the modeling in [12], we take $\varepsilon=2$ and $p_{0}=p_{1}=1/2$, so that $\mathcal{D}(\varepsilon\varrho)\circ\pi_{1}^{-1}=\operatorname{Beta}(1,1)$, where $\pi_{1}:M_{1}\to[0,1]$ is the projection, $\nu\mapsto\nu(\{1\})$. We will use and compare two different values of $\kappa$ (which is denoted by $c$ in [12]). For the data, we have $M=320$ and $N_{m}=9$ for all $m$. Our row counts, $\overline{y}$, are given by $\overline{y}_{m1}=r_{m}$ and $\overline{y}_{m0}=9-r_{m}$.
We first consider $\kappa=1$. As in [12], we generated $K=10000$ weighted simulations. In this case, our effective sample size was approximately 244. (For comparison, in [12], Liu reported an effective sample size of 227 for the case $\kappa=1$.) The unknown probability of success for a new thumbtack is given by $\theta_{321,1}$, and (6.6) gives
$\mathcal{L}(\theta_{321,1}\mid Y=y)\approx\frac{1}{321}\bigg{(}\operatorname{Beta}(1,1)+\sum_{m=1}^{320}\frac{\sum_{k=1}^{10000}V_{k}\delta(\theta_{m,1}^{*,k})}{\sum_{k=1}^{10000}V_{k}}\bigg{)}.$
An approximate density for this measure is given in Figure 2(a).
We next consider $\kappa=10$, again using $K=10000$, which generated an effective sample size of about 388 (compared to 300 in [12] for the same value of $\kappa$). This time, using (6.6) gives
$\mathcal{L}(\theta_{321,1}\mid Y=y)\approx\frac{1}{330}\bigg{(}10\operatorname{Beta}(1,1)+\sum_{m=1}^{320}\frac{\sum_{k=1}^{10000}V_{k}\delta(\theta_{m,1}^{*,k})}{\sum_{k=1}^{10000}V_{k}}\bigg{)}.$
Note that in this second case, the simulated values $\theta_{m,1}^{*,k}$ and their corresponding weights $V_{k}$ were all regenerated. An approximate density for this measure is given in Figure 2(b).
(a) $\kappa=1$
(b) $\kappa=10$
Figure 2: Approximate density of $\mathcal{L}(\theta_{321,1}\mid Y=y)$
As in the previous example, these approximate densities were constructed using Gaussian kernel density estimation. Their respective bandwidths are $h\approx 0.105$ and $h\approx 0.096$. The graphs in Figure 2 are qualitatively similar to their counterparts in [12], but with minor differences. It is difficult, though, to make a direct comparison. Although Gaussian kernel smoothing was also used in [12], details about the smoothing were not provided. For instance, the bandwidths used to produce the graphs in [12] were not reported therein.
6.4 Amazon reviews
The model in [12] only covers agents with two possible actions, such as coins and thumbtacks. The IDP, though, can handle agents whose range of possible actions is arbitrary.
Imagine, then, that we discover a seller on Amazon that has 50 products. Their products have an average rating of 2.4 stars out of 5. Some products have almost 100 ratings, while others have only a few. On average, the products have 23 ratings each. In this case, the agents are the products and the actions are the ratings that each product earns. Each individual rating must be a whole number of stars between 1 and 5, inclusive. Hence, each action has 5 possible outcomes. The data used for this hypothetical seller is given in Table LABEL:Tb:reviews.
Table 2: Reviews for 50 different products from a given seller
product # 1 star 2 stars 3 stars 4 stars 5 stars # reviews average
1 9 25 15 41 0 90 2.98
2 21 28 18 1 3 71 2.11
3 16 11 21 11 0 59 2.46
4 3 9 37 0 3 52 2.83
5 11 0 36 0 5 52 2.77
6 16 16 4 15 0 51 2.35
7 30 3 15 0 0 48 1.69
8 12 9 17 1 7 46 2.61
9 13 13 18 1 0 45 2.16
10 23 2 0 14 0 39 2.13
11 11 4 6 7 10 38 3.03
12 6 3 21 0 5 35 2.86
13 14 9 0 5 2 30 2.07
14 4 25 0 0 0 29 1.86
15 8 7 2 10 0 27 2.52
16 5 4 6 10 0 25 2.84
17 6 10 9 0 0 25 2.12
18 11 1 2 3 7 24 2.75
19 20 3 0 0 0 23 1.13
20 6 9 4 2 1 22 2.23
21 5 1 3 8 1 18 2.94
22 9 1 5 2 1 18 2.17
23 5 7 3 1 1 17 2.18
24 0 3 12 0 2 17 3.06
25 1 11 1 3 1 17 2.53
26 0 3 0 6 7 16 4.06
27 2 2 8 3 1 16 2.94
28 6 5 1 3 0 15 2.07
29 6 6 1 2 0 15 1.93
30 0 8 2 4 0 14 2.71
31 8 5 1 0 0 14 1.5
32 5 0 8 0 1 14 2.43
33 0 0 13 0 0 13 3
34 5 4 1 2 0 12 2
35 6 2 0 3 0 11 2
36 4 7 0 0 0 11 1.64
37 0 1 6 4 0 11 3.27
38 5 5 0 1 0 11 1.73
39 5 6 0 0 0 11 1.55
40 1 2 2 4 1 10 3.2
41 4 1 3 1 0 9 2.11
42 4 1 1 0 0 6 1.5
43 3 1 0 1 0 5 1.8
44 3 0 1 0 0 4 1.5
45 1 2 0 0 0 3 1.67
46 0 1 2 0 0 3 2.67
47 2 1 0 0 0 3 1.33
48 0 0 2 0 0 2 3
49 0 1 0 1 0 2 3
50 0 0 1 1 0 2 3.5
To model this example we take $L=5$, so that $S=\{0,1,2,3,4\}$, where $\ell\in S$ represents an $(\ell+1)$-star review. We take $\kappa=10$, $\varepsilon=5$, and $p_{\ell}=1/5$ for each $\ell\in S$. For the data, we have $M=50$ and the number $N_{m}$ is the total number of reviews given to the $m$th product. For our row counts, the number $\overline{y}_{m\ell}$ is the total number of $(\ell+1)$-star reviews given to the $m$th product. In this example, we generated $K=100000$ weighted simulations, and obtained an effective sample size of about 561. In computing the simulation weights as in (6.4), we used a log scale factor of 28.8.
As with the pressed penny machine, we begin by considering a hypothetical new product from this seller. The quality of this 51st product can be characterized by the vector $\theta_{51}=(\theta_{51,0},\theta_{51,1},\theta_{51,2},\theta_{51,3},\theta_{51,4})$, since $\theta_{51,\ell}$ is the (unknown) probability that the product will receive an $(\ell+1)$-star review. The long-term average rating of this product over many reviews will be $A(\theta_{51})$, where $A(x)=\sum_{\ell=0}^{4}\ell x_{\ell}$. According to (6.6), we have
$\mathcal{L}(A(\theta_{51})\mid Y=y)\approx\frac{1}{60}\left({10\operatorname{Dir}(1,1,1,1,1)\circ A^{-1}+\sum_{m=1}^{50}\frac{\sum_{k=1}^{100000}V_{k}\delta(A(\theta_{m}^{*,k}))}{\sum_{k=1}^{100000}V_{k}}}\right).$
Using this, we have $E[A(\theta_{51})\mid Y=y]\approx 2.54$, meaning that the expected long-term average rating of a new product is a little more than 2.5. For a more informative look at the quality of a new product, an approximate density for $\mathcal{L}(A(\theta_{51})\mid Y=y)$ is given in Figure 3(a).
(a) $m=51$, mean: 2.54
(b) $m=50$, mean: 2.83
(c) $m=26$, mean: 3.8
Figure 3: Approximate densities for $\mathcal{L}(A(\theta_{m})\mid Y=y)$
This graph in Figure 3(a) shows a bimodal distribution, meaning that we can expect the average ratings of future products to cluster around 2 and 3 stars.
After having considered a hypothetical new product, we turn our attention to the 50 products that have already received reviews. Consider, for instance, the 50th product. This product has a 3.5-star average rating, but only 2 reviews. To see the effect of these 2 reviews on the expected long-term rating, we apply (6.5) with $\Phi((x_{m\ell}))=A(x_{50})$ to obtain
$\mathcal{L}(A(\theta_{50})\mid Y=y)\approx\frac{\sum_{k=1}^{100000}V_{k}\delta(A(\theta_{50}^{*,k}))}{\sum_{k=1}^{100000}V_{k}}.$
This gives $E[A(\theta_{50})\mid Y=y]\approx 2.83$, and an approximate density for $\mathcal{L}(A(\theta_{50})\mid Y=y)$ is given in Figure 3(b). According to the model, the 50th product’s two reviews (a 3-star and a 4-star review) have transformed the graph in Figure 3(a) to the graph in 3(b), and increased its expected long-term average rating from 2.54 to 2.83.
We can similarly look at the 26th product. This product has an average rating of 4.06, but it only has 16 reviews. Using (6.5) as above, we obtain $E[A(\theta_{26})\mid Y=y]\approx 3.8$, and an approximate density for $\mathcal{L}(A(\theta_{26})\mid Y=y)$ as given in Figure 3(c).
6.5 The gamer distribution
In our previous examples, we took $\bm{p}$ to be a uniform measure. That is, we took $p_{\ell}=1/L$ for all $\ell$. Our final example will be presented in Section 6.6 and is concerned with video game leaderboards. In that example, to have plausible results that match our intuition about video games, it will not be sufficient to let $\bm{p}$ be uniform. Instead, we will construct $\bm{p}$ from the continuous distribution described in this section.
Let $r$, $c$, and $\alpha$ be positive real numbers. A nonnegative random variable $X$ is said to have the gamer distribution with parameters $r$, $c$, and $\alpha$, denoted by $X\sim\operatorname{Gamer}(r,c,\alpha)$, if $X$ has density
[…]
for $x>0$. The fact that this is a probability density function is shown in Proposition 6.3 below. Note that if $X\sim\operatorname{Gamer}(r,c,\alpha)$ and $s>0$, then $sX\sim\operatorname{Gamer}(r,sc,\alpha)$.
The gamer distribution is meant to model the score of a random player in a particular single-player game. The game is assumed to have a structure in which the player engages in a sequence of activities that can result in success or failure. Successes increase the player’s score. Failures bring the player closer to a termination event, which causes the game to end.
The distribution of scores at the higher end of the player skill spectrum has a power law decay with exponent $r$. More specifically, there is constant $K$ such that $P(X>x)\approx Kx^{-r}$ for large values of $x$. For small values of $x$, the distribution of $X$ looks like a gamma distribution.
The parameter $c$ indicates the average score of players at the lower end of the skill spectrum, which make up the bulk of the player base. The parameter $\alpha$ is connected to the structure of the game. Higher values of $\alpha$ indicate a more forgiving game in which the termination event is harder to trigger. See below for more on the meaning of these parameters.
The gamer distribution can be seen as a mixture of gamma distributions, where the mixing distribution is Pareto. More specifically, it is straightforward to prove the following.
Proposition 6.3.
Let $r,c>0$ and let $M$ have a Pareto distribution with minimum value $c$ and tail index $r$. That is, $P(M>m)=(m/c)^{-r}$ for $m>c$. If $X\mid M\sim\operatorname{Gamma}(\alpha,\alpha/M)$, then $X\sim\operatorname{Gamer}(r,c,\alpha)$.
[…]
$f(x)=\frac{rc^{r}}{\alpha^{r}\,\Gamma(\alpha)}\,x^{-r-1}\,\gamma\left(\alpha+r,\frac{\alpha x}{c}\right).$ (6.8)
Since $\gamma(\beta,u)\to\Gamma(\beta)$ as $u\to\infty$, we have
$f(x)\sim\frac{\Gamma(\alpha+r)}{\alpha^{r}\,\Gamma(\alpha)}\,\frac{rc^{r}}{x^{r+1}}$
as $x\to\infty$. In other words, the density of $X$ is asymptotically proportional to the density of $M$ as $x\to\infty$.
For small values of $x$, note that $\gamma(\beta,u)\sim u^{\beta}e^{-u}$ as $u\to 0$. Hence, if we introduce the parameter $\lambda=\alpha/c$, then
$f(x)\sim\frac{r}{\lambda^{r}\Gamma(\alpha)}x^{-r-1}(\lambda x)^{\alpha+r}e^{-\lambda x}=r\,\frac{\lambda^{\alpha}}{\Gamma(\alpha)}\,x^{\alpha-1}e^{-\lambda x}$
as $x\to 0$. In other words, the density of $X$ is asymptotically proportional to the density of $\operatorname{Gamma}(\alpha,\alpha/c)$ as $x\to 0$. Since $\operatorname{Gamma}(\alpha,\alpha/c)$ has mean $c$, the parameter $c$ can be understood as the average score of players at the lower end of the skill spectrum.
To understand $\alpha$, we look to the fact that $X\mid M\sim\operatorname{Gamma}(\alpha,\alpha/M)$. Given $M$, we can think of $X$ as being driven by $\alpha$ exponential clocks, each with mean $M/\alpha$. Each clock represents a time to failure, and when all clocks expire, the player has reached the termination event. Since $\alpha$ denotes the number of such clocks, a higher value of $\alpha$ indicates that more failures are needed to trigger the end of the game. We also have $\operatorname{Var}(X\mid M)=M^{2}/\alpha$. Hence, $\alpha$ can also be understood through the fact that $1/\sqrt{\alpha}$ is the coefficient of variation of $X$ given $M$.
For computational purposes, it may be more efficient to rewrite (6.8) in terms of the logarithm of the gamma function and the regularized lower incomplete gamma function, $P(\beta,u)=\gamma(\beta,u)/\Gamma(\beta)$. In this case, we have
$f(x)=r\left(\frac{c}{\alpha}\right)^{r}\exp(\log\Gamma(\alpha+r)-\log\Gamma(\alpha))x^{-r-1}P\left(\alpha+r,\frac{\alpha x}{c}\right)$
for $x>0$.
6.6 Video game leaderboards
For our final example, we consider a single-player video game in which an individual player tries to score as many points as possible before the game ends. If $X$ is a random score of a random player, then we will assume that $X\sim\operatorname{Gamer}(r,c,\alpha)$, where $r=7/3$, $c=28$, and $\alpha=3$. The values of these parameters are arbitrarily chosen for the sake of the example. The values of $r$ and $c$ give the distribution a mean of about 50 and a decay rate that approximately matches the decay rate in the global Tetris leaderboard (see https://kirjava.xyz/tetris-leaderboard/). The choice $\alpha=3$ indicates a game in which the player has 3 “lives,” which is a typical gaming structure, especially in classic arcade video games. Finally, we assume that the actual score displayed by the game is rounded to the nearest integer and capped at 499.
A group of 10 friends get together and play this game. Each friend plays the game a different number of times. In this case, the agents are the players and the actions are the scores they earn each time they play.
The 10 friends all have their own usernames that they use when playing the game. The usernames are Asparagus Soda, Goat Radish, Potato Log, Pumpkins, Running Stardust, Sweet Rolls, The Matrix, The Pianist Spider, The Thing, and Vertigo Gal. We will consider three different scenarios for this example.
6.6.1 Players with matching scores
In our first scenario, the 10 friends generate the scores given in Table 3. Note that in that table, the scores are listed in increasing order. To get an overview of the data, we can place the 10 players in a leaderboard, ranked by their high score, as shown in Table 4.
Username Scores
Pumpkins 12, 21, 25, 25, 26, 27, 30, 33, 34, 34, 36, 42, 44, 44, 48, 55, 67, 69
Potato Log 18, 21, 21, 22, 23, 25, 29, 29, 32, 33, 47, 53, 54, 56, 57, 65, 75
The Thing 10, 16, 16, 19, 19, 25, 25, 26, 29, 32, 35, 37, 42, 44, 59, 60
Running Stardust 23, 38, 62, 71, 138, 149, 151
Sweet Rolls 15, 23, 56, 71, 98, 130
Vertigo Gal 10, 30, 40, 56, 87, 92
Asparagus Soda 17, 43, 55
The Matrix 11, 15
Goat Radish 38
The Pianist Spider 3
Table 3: Player scores for Video Game Scenario 1
rank name hi score avg score IDP avg # games
1 Running Stardust 151 90 80 7
2 Sweet Rolls 130 66 55 6
3 Vertigo Gal 92 52 52 6
4 Potato Log 75 39 39 17
5 Pumpkins 69 37 38 18
6 The Thing 60 31 32 16
7 Asparagus Soda 55 38 40 3
8 Goat Radish 38 38 71 1
9 The Pianist Spider 32 32 37 1
10 The Matrix 15 13 43 2
Table 4: Leaderboard for Video Game Scenario 1
To model this scenario, we take $L=500$, so that $S=\{0,1,\ldots,499\}$. We take $\kappa=\varepsilon=1$ and let
$p_{\ell}=\begin{cases}F(\ell+0.5)-F(\ell-0.5)&\text{if $0\leq\ell<499$},\\ 1-F(498.5)&\text{if $\ell=499$},\end{cases}$
where $F$ is the distribution function of a $\operatorname{Gamer}(7/3,28,3)$ distribution. For the data, we have $M=10$, the number $N_{m}$ is the number of scores in the $m$th row of Table 3, and $y_{mn}$ is the $n$th score in the $m$th row. Note that since the model only depends on $y_{mn}$ through the row counts $\overline{y}_{m\ell}$, the order in which the scores are listed in the vector $y_{m}$ is not relevant. In this scenario, we generated $K=40000$ weighted simulations, and obtained an effective sample size of about 326. In computing the simulation weights as in (6.4), we used a log scale factor of 42.
The long-term average score of the player in the $m$th row of Table 3 will be $A(\theta_{m})$, where $A(x)=\sum_{\ell=0}^{499}\ell x_{\ell}$. For example, using (6.5), the expected long-term average score of Running Stardust is $E[A(\theta_{4})\mid Y=y]\approx 79.65$. These conditional expectations, rounded to the nearest integer, are shown in the “IDP avg” column of Table 4.
Looking at these averages, we can see at least two players whose numbers seem unusual. The first is Goat Radish. They played only one game and scored a 38, which is a relatively low score compared to the rest of the group. And yet the IDP model has given them an expected long-term average score of 71. Not only is this counterintuitive, it is also inconsistent with how the model treated The Pianist Spider.
The reason for this behavior can be seen in Table 3. There is only one other player that managed to score exactly 38 in one of their games: Running Stardust. So from the model’s perspective, there is a reasonable chance that Goat Radish and Running Stardust have similar scoring tendencies. Since Running Stardust happens to be the top player, this leads to an unusually high long-term estimate for Goat Radish.
Our intuition is able to dismiss this line of reasoning because we know, for instance, that there is very little difference between a score of 38 and 39. Had Goat Radish scored a 39 instead, our predictions should not change that much. But we only know this because we are viewing the positive real numbers as more than just a set. We are viewing them as a totally ordered set with the Euclidean metric. The IDP model is not designed to utilize these properties of the state space. From its perspective, the number “38” is just a label. It is nothing more than the name of a particular element of the state space, and it happens to be an element that only two players were able to hit.
We see similar behavior in the model’s forecast for The Matrix, who scored an 11 and a 15 in their two games. No one else scored an 11, but exactly one other player managed to score exactly 15, and that was Sweet Rolls, who happens to be the second best player. Just as with Goat Radish, this causes the model to generate an unintuitively high value for The Matrix’s long-term average score.
To test this explanation, we are led to our second scenario.
6.6.2 Matching scores removed
The scores in our second scenario are the same as in our first, but we changed Goat Radish’s 38 to a 39, and The Matrix’s 15 to a 14. (See Table 5.) The scores 14 and 39 are unique in that no other player achieved exactly those scores. We reran the model, again generating $K=40000$ weighted simulations. This time, we obtained an effective sample size of about 22.3. To save time, we deleted the two heaviest simulations, leaving $K=39998$ simulations with an effective sample size of about 1099. The new expected long-term averages are shown in Table 6.
We now see that Goat Radish and The Matrix have lower, more reasonable long-term averages according to the model. Likewise, Running Stardust and Sweet Rolls have slightly higher averages. In the first scenario, their averages were brought down because of their associations with Goat Radish and The Matrix.
Username Scores
Pumpkins 12, 21, 25, 25, 26, 27, 30, 33, 34, 34, 36, 42, 44, 44, 48, 55, 67, 69
Potato Log 18, 21, 21, 22, 23, 25, 29, 29, 32, 33, 47, 53, 54, 56, 57, 65, 75
The Thing 10, 16, 16, 19, 19, 25, 25, 26, 29, 32, 35, 37, 42, 44, 59, 60
Running Stardust 23, 38, 62, 71, 138, 149, 151
Sweet Rolls 15, 23, 56, 71, 98, 130
Vertigo Gal 10, 30, 40, 56, 87, 92
Asparagus Soda 17, 43, 55
The Matrix 11, 14
Goat Radish 39
The Pianist Spider 3
Table 5: Player scores for Video Game Scenario 2
rank name hi score avg score IDP avg # games
1 Running Stardust 151 90 84 7
2 Sweet Rolls 130 66 62 6
3 Vertigo Gal 92 52 51 6
4 Potato Log 75 39 39 17
5 Pumpkins 69 37 38 18
6 The Thing 60 31 31 16
7 Asparagus Soda 55 38 39 3
8 Goat Radish 38 38 43 1
9 The Pianist Spider 32 32 37 1
10 The Matrix 15 13 28 2
Table 6: Leaderboard for Video Game Scenario 2
6.6.3 Players with only a few games
In our third scenario, the ten friends generated the scores in Table 7. We use the same $L$, $\kappa$, $\varepsilon$, $\bm{p}$, and $M$ as in the first scenario. Also as it was there, the number $N_{m}$ is the number of scores in the $m$th row of Table 7, and $y_{mn}$ is the $n$th score in the $m$th row. Note, however, that the username in the $m$th row has changed in the current scenario. We again used a log scale factor of 42 and generated $K=40000$ weighted simulations, obtaining an effective sample size of about 39. This time, we deleted the 26 heaviest simulations, leaving $K=39974$ simulations and an effective sample size of about 207. As before, the resulting long-term expected averages, $E[A(\theta_{m})\mid Y=y]$, are shown in Table 8.
In this example, we focus our attention on Asparagus Soda, who it situated at No. 4 on the leaderboard, but played the game only once. The question is, does he deserve to be at No. 4? Is he truly the fourth-best player among the ten friends? For example, Potato Log, who is at No. 3, played the game 20 times and only managed to get a high score of 87. Asparagus Soda almost matched that high score in a single attempt. Intuitively, it seems clear that Asparagus Soda is the better player and should rank higher than Potato Log.
It is less clear how Asparagus Soda compares to Pumpkins, the No. 2 player. Neither of them made a lot of attempts, but Asparagus Soda has the higher average score. Which one is more likely to have the higher long-term average score? If they had a contest where they each played a single game and the higher score wins, who should we bet on?
Username Scores
Vertigo Gal 45, 100, 118, 121, 125, 130, 133, 145, 161, 173, 173, 187, 190, 192, 193, 200, 220, 223, 256, 275, 314, 354, 388, 475, 524
Potato Log 4, 13, 13, 16, 19, 19, 19, 19, 23, 24, 25, 26, 31, 38, 41, 43, 44, 47, 51, 87
The Thing 4, 6, 9, 19, 25, 27, 28, 38, 39, 40
The Matrix 13, 15, 17, 32, 32, 61, 78
Running Stardust 21, 23, 51, 61, 65
Goat Radish 23, 25, 34, 51
Pumpkins 49, 65, 84, 117
Sweet Rolls 26, 65
Asparagus Soda 86
The Pianist Spider 62
Table 7: Player scores for Video Game Scenario 3
rank name hi score avg score IDP avg # games
1 Vertigo Gal 475 207 198 25
2 Pumpkins 117 79 72 4
3 Potato Log 87 30 31 20
4 Asparagus Soda 86 86 67 1
5 The Matrix 78 35 37 7
6 Running Stardust 65 44 45 5
6 Sweet Rolls 65 46 52 2
8 The Pianist Spider 62 62 56 1
9 Goat Radish 51 33 34 4
10 The Thing 40 24 26 10
Table 8: Leaderboard for Video Game Scenario 3
Asparagus Soda vs. Potato Log.
Looking at Table 7, we see that Asparagus Soda corresponds to $m=9$ and Potato Log corresponds to $m=2$. Table 8 shows us that $E[A(\theta_{9})\mid Y=y]\approx 67$ and $E[A(\theta_{2})\mid Y=y]\approx 31$. In other words, the IDP model gives Asparagus Soda a much higher expected long-term average score than Potato Log. This confirms our intuition that Asparagus Soda is the better player. But because Asparagus Soda played only one game, the model should have a lot more uncertainty surrounding Asparagus Soda’s forecasted mean. To see this, we can compare approximate densities for $\mathcal{L}(A(\theta_{9})\mid Y=y)$ and $\mathcal{L}(A(\theta_{2})\mid Y=y)$. (See Figure 4.)
(a) density for $\mathcal{L}(A(\theta_{9})\mid Y=y)$
(b) density for $\mathcal{L}(A(\theta_{2})\mid Y=y)$
(c) density for $\mathcal{L}(C(\theta_{9},\theta_{2})\mid Y=y)$
Figure 4: Asparagus Soda ($m=9$) vs. Potato Log ($m=2$)
As is visually evident, Asparagus Soda’s density is supported on a much wider interval. In this way, the model acknowledges the possibility that Asparagus Soda’s actual long-term average score is lower than Potato Log’s. The probability that this is the case is $P(A(\theta_{9})<A(\theta_{2})\mid Y=y)$. If we define $\Phi:\mathbb{R}^{10\times 500}\to\mathbb{R}$ by $\Phi((x_{m\ell}))=A(x_{9})-A(x_{1})$, then we can use (6.5) to obtain $P(A(\theta_{9})<A(\theta_{2})\mid Y=y)\approx 0.049$. In other words, according to the model, there is a 95% chance that Asparagus Soda is a better player than Potato Log.
Now suppose the two of them had a contest in which they each played the game once and the higher score wins. What is the probability that Asparagus Soda would win this contest? If we define $C:\mathbb{R}^{500}\times\mathbb{R}^{500}\to\mathbb{R}$ by $C(x,y)=\sum_{\ell>\ell^{\prime}}x_{\ell}y_{\ell^{\prime}}$ and then $C(\theta_{9},\theta_{2})$ is the (unknown) probability that Asparagus Soda beats Potato Log in this single-game contest. The actual probability, given our observations $Y=y$, is then $E[C(\theta_{9},\theta_{2})\mid Y=y]$, which, according to (6.5), is approximately $0.786$. That is, Asparagus Soda has about a 79% chance of beating Potato Log in a contest involving a single play of the game. To visualize the uncertainty around this probability, we can graph an approximate density for $\mathcal{L}(C(\theta_{9},\theta_{2})\mid Y=y)$. This is done in Figure 4(c). The graph shows that although the conditional mean of $C(\theta_{9},\theta_{2})$ is about 79%, the conditional mode is much higher.
Asparagus Soda vs. Pumpkins.
We now turn our attention to comparing Asparagus Soda, who played only once, to Pumpkins, who played four times. (See Figure 5.)
(a) density for $\mathcal{L}(A(\theta_{9})\mid Y=y)$
(b) density for $\mathcal{L}(A(\theta_{7})\mid Y=y)$
(c) density for $\mathcal{L}(C(\theta_{9},\theta_{7})\mid Y=y)$
Figure 5: Asparagus Soda ($m=9$) vs. Pumpkins ($m=7$)
Looking at Table 7, we see that Asparagus Soda corresponds to $m=9$ and Pumpkins corresponds to $m=7$. Table 8 shows us that $E[A(\theta_{9})\mid Y=y]\approx 67$ and $E[A(\theta_{7})\mid Y=y]\approx 72$. We can visualize the model’s uncertainty around Pumpkins’ expected long-term average by graphing an approximate density for $\mathcal{L}(A(\theta_{7})\mid Y=y)$. This is done in Figure 5(b).
Visually comparing this graph with the corresponding one for Asparagus Soda in Figure 5(a), we see that the two long-term averages have comparable degrees of uncertainty. Using (6.5), we have $P(A(\theta_{9})<A(\theta_{2})\mid Y=y)\approx 0.625$, meaning there is a 62% chance that Pumpkins is the better player.
We can also consider a single-game contest between Asparagus Soda and Pumpkins. As above, we can use (6.5) to compute $E[C(\theta_{9},\theta_{7})\mid Y=y]\approx 0.484$, meaning that Asparagus Soda has a 48% chance of beating Pumpkins in a single-game contest. To visualize the uncertainty around this probability, we can graph an approximate density for $\mathcal{L}(C(\theta_{9},\theta_{7})\mid Y=y)$. (See Figure 5(c).)